COLOCALIZING SUBCATEGORIES AND COSUPPORT

DAVE BENSON, SRIKANTH B. IYENGAR, AND HENNING KRAUSE

Abstract. The Hom closed colocalizing subcategories of the stable module category of a finite group are classified. Along the way, the colocalizing subcategories of the homotopy category of injectives over an exterior algebra, and the derived category of a formal commutative differential graded algebra, are classified. To this end, and with an eye towards future applications, a notion of local homology and cosupport for triangulated categories is developed, building on earlier work of the authors on local cohomology and support.

Contents

1. Introduction  1
2. (Co)localization functors on triangulated categories  4
3. Local cohomology and support  6
4. Local homology and cosupport  8
5. 𝔭-local and 𝔭-complete objects  13
6. Axioms for support and cosupport  15
7. Change of rings and categories  17
8. Tensor triangulated categories  22
9. Costratification  24
10. Formal dg algebras  27
11. Finite groups  28
Appendix A. Localization functors and their adjoints  34
References  37

1. Introduction

Let G be a finite group and k a field of characteristic p, dividing the order of G, and StMod(kG) the stable module category of possibly infinite dimensional kG-modules. We write V_{G} for the set of all homogeneous prime ideals except the maximal ideal in the cohomology algebra H^{∗}(G, k) of G, and V_{G}(M) for the support of any M ∈ StMod(kG), defined by Benson, Carlson and Rickard [3] when k is algebraically closed, and extended in [5] to all fields. One of the main results in this work is a classification of the colocalizing subcategories of StMod(kG):

Theorem 1.1. The map that assigns to each subset U ⊆ V_{G} the subcategory

\{N ∈ StMod(kG) | \underline{Hom}_{kG}(M, N) = 0 for all M with V_{G}(M) ⊆ U\}

2010 Mathematics Subject Classification. 20J06(primary); 13D45, 16E45, 18E30.

Key words and phrases. colocalizing subcategory, cosupport, local homology, localizing subcategory, stable module category, triangulated category.

The research of the first and second authors was undertaken during visits to the University of Paderborn, each supported by a research prize from the Humboldt Foundation. The research of the second author was also partly supported by NSF grant DMS 0903493.

gives a bijection between subsets of V_{G} and colocalizing subcategories of StMod(kG) that are closed under tensor product with simple kG-modules.

A colocalizing subcategory C is by definition a full triangulated subcategory that is closed under set-indexed products. Such a subcategory is closed under tensor product with simples if and only if it is Hom closed: If N is in C, so is Hom_{k}(M, N) for any M ∈ StMod(kG). Theorem 1.1 complements the classification of the localizing subcategories of StMod(kG) from [5, Theorem 10.3]. Combining them gives a remarkable bijection:

Corollary 1.2. The map sending a localizing subcategory S of StMod(kG) to S^{⊥} induces a bijection

\{ \begin{matrix}tensor closed localizing \\ subcategories of StMod(kG) \end{matrix} \} [figure: ∼]\{ \begin{matrix}Hom closed colocalizing \\ subcategories of StMod(kG) \end{matrix}\} .

The inverse map sends a colocalizing subcategory S to ^{⊥}S.

Theorem 1.1 and Corollary 1.2, proved in Section 11, are analogues of recent results of Neeman [21] on the derived category of a noetherian commutative ring.

The definition of the inverse of the map in Theorem 1.1 involves a notion of cosupport for a module M in StMod(kG), introduced in this work to be the subset

cosupp_{G} M = {𝔭 ∈ V_{G} | Hom_{k}(κ_{𝔭}, M) is not projective},

with κ_{𝔭} the Rickard idempotent module associated to 𝔭, constructed in [3]. Recall that the support of M is {𝔭 ∈ V_{G} | κ_{𝔭} ⊗ M is not projective}. The inverse map in Theorem 1.1 assigns to a subcategory C of StMod(kG) the complement in V_{G} of the set ⋃_{M∈C} cosupp_{G} M.

The proof of Theorem 1.1 is modelled on that of [5, Theorem 10.3], where localizing subcategories of StMod(kG) are classified. It involves a sequence of changes of category, for which reason it has been necessary to develop a theory of cosupport for objects in triangulated categories, along the lines for the one for support in our earlier work [4, 5, 6].

The context is a compactly generated triangulated category T with set-indexed coproducts endowed with an action of a graded commutative noetherian ring R; meaning, a homomorphism R → Z^{∗}(T) of graded rings from R to the graded center of T. For each 𝔭 in Spec R, the set of homogeneous prime ideals in R, we introduce a local homology functor Λ^{𝔭}, constructed as a right adjoint to the local cohomology functor, Γ_{𝔭}, from [4], and define the cosupport of an object X of T by

cosupp_{R} X = {𝔭 ∈ Spec R | Λ^{𝔭}X ≠ 0} .

In the first part of this paper, Sections 2 to 5, we establish salient properties of local homology and cosupport; for instance, that the maximal elements with respect to inclusion in the cosupport and the support of any object X in T coincide:

max(cosupp_{R} X) = max(supp_{R} X) .

This is proved as part of Theorem 4.13. It follows that cosupp_{R} X = ∅ if and only if supp_{R} X = ∅, which is equivalent to X = 0 by [4, Theorem 5.2]. These results suggest a close connection between the support and cosupport. However, while the support of an object is well-understood, the cosupport remains a mysterious entity. For instance, the only complete results we could obtain for finitely generated modules over commutative noetherian rings are given in Propositions 4.18 and 4.19.

From Section 8 onwards we turn to colocalizing subcategories of T, focusing on the case when T is tensor triangulated with a canonical R-action, meaning an action induced by a homomorphism R → End_{T}^{∗}(𝟙), where 𝟙 is the unit for the tensor product on T. This is the context of the main results of this work, and the rest of this introduction. The category T admits an internal function object, denoted Hom(X, Y), and it is natural to examine the Hom closed colocalizing subcategories of T. A useful result concerning these is that for each X ∈ T there is an equality

Coloc_{T}^{Hom}(X) = Coloc_{T}^{Hom}({Λ^{𝔭}X | 𝔭 ∈ Spec R})

which is a form of local-global principle for colocalizing subcategories. This statement is Theorem 8.6 and an analogue of such a local-global principle for localizing subcategories in [5, Theorem 3.6]. The theorem is a first step in our approach to the problem of classifying the Hom closed colocalizing subcategories of T, for it permits one to reduce it to the classification problem for Λ^{𝔭}T, the essential image of the functor Λ^{𝔭}, for each 𝔭 ∈ Spec R; see Proposition 9.1. We note that Λ^{𝔭}T is itself colocalizing and Hom closed; see Propositions 4.16 and 8.3.

The following definition thus naturally emerges: the tensor triangulated category T is costratified by R if for each 𝔭 ∈ Spec R there are no non-trivial Hom closed colocalizing subcategories in Λ^{𝔭}T. Given the discussion above, it is clear that when this property holds the map assigning to a subcategory C the subset ⋃_{X∈C} cosupp_{R} X of Spec R sets up a bijection

\{ \begin{matrix}Hom closed colocalizing \\ subcategories of T \end{matrix} \} ⟶^{∼} \{ subsets of supp_{R} T \} .

This bijection is Corollary 9.2 and was the main reason for our interest in the costratification condition. However, there are other remarkable consequences that follow from it. For instance, we prove in Theorem 9.7: if T is costratified by R, it is also stratified by R, meaning that there are no proper tensor closed localizing subcategories of Γ_{𝔭}T; see [5, 6]. One consequence is that if T is costratified by R then there is a bijection, analogous to the one in Corollary 1.2, between the tensor closed localizing subcategories and the Hom closed colocalizing subcategories of T, via left and right perp; see Corollary 9.9.

In Theorem 9.5 we prove that if T is stratified by R there is an equality

cosupp_{R} Hom(X, Y) = supp_{R} X ∩ cosupp_{R} Y for all X, Y ∈ T.

This is an analogue of the tensor product theorem for support [6, Theorem 7.3]. It follows that one gets

Hom_{T}^{∗}(X, Y) = 0 ⟺ supp_{R} X ∩ cosupp_{R} Y = ∅

provided that the tensor identity generates T; see Corollary 9.6. This is a surprisingly complete result, for it is often difficult to obtain precise conditions under which there are non-zero maps between objects in a triangulated category.

In Section 11 we prove Theorem 1.1, by establishing that the tensor triangulated category StMod(kG) is costratified by the canonical action of H^{∗}(G, k). Along the way we prove that the following tensor triangulated categories are costratified:

• The derived category of a formal dg algebra whose cohomology is graded commutative and noetherian; see Theorem 10.3.

• The homotopy category of graded injectives over an exterior algebra, viewed as a dg algebra with zero differential; see Theorem 10.4.

• The homotopy category of complexes of injective kG-modules, where G is

a finite group; see Theorem 11.10. The proofs of these results use much of the material on cosupport and local homology in the preceding sections, as well as results on their behavior under changes of categories, studied in Section 7. Specialized to the case of a commutative noetherian ring, viewed as a dg algebra concentrated in degree 0, the first item in the preceding list is Neeman’s theorem, mentioned at the beginning, which was the inspiration for the results described in this article.

Acknowledgments. It is a pleasure to thank Amnon Neeman for carefully reading a preliminary version of this manuscript, and the referee for suggestions regarding the presentation of this material.

2. (Co)localization functors on triangulated categories

In this section we collect basic facts about localization and colocalization functors on triangulated categories required in this work; see [4, §3] for details.

Let T be a triangulated category which admits set-indexed products and coproducts. We write Σ for the suspension functor on T. The kernel of an exact functor F: T → T is the full subcategory

Ker L = {X ∈ T | LX = 0} ,

while the essential image of F is the full subcategory

Im F = {X ∈ T | X ≅ F Y for some Y in T}.

A localizing subcategory of T is a full triangulated subcategory that is closed under taking all coproducts. We write Loc_{T}(C) for the smallest localizing subcategory containing a given class of objects C in T, and call it the localizing subcategory generated by C. Analogously, a colocalizing subcategory of T is a full triangulated subcategory that is closed under taking all products, and Coloc_{T}(C) denotes the colocalizing subcategory of T that is cogenerated by C.

A localization functor L: T → T is an exact functor that admits for each X in T a natural morphism ηX: X → LX, called adjunction, such that L(ηX) is an isomorphism and L(ηX) = η(LX). A functor Γ: T → T is a colocalization functor if its opposite functor Γ^{op} : T^{op} → T^{op} is a localization functor; the corresponding natural morphism θX: Γ X → X is called coadjunction.

A localization functor L: T → T is essentially determined by its kernel, which is a localizing subcategory of T, for it coincides with the kernel of a functor that admits a right adjoint; see [4, Lemma 3.1]. The natural transformation η: Id_{T} → L induces for each object X in T a natural exact localization triangle

Γ X ⟶ X ⟶ LX ⟶

This exact triangle gives rise to an exact functor Γ: T → T with

Ker L = Im Γ and Ker Γ = Im L.

The functor Γ is a colocalization, and each colocalization functor on T arises in this way. This yields a natural bijection between localization and colocalization functors on T. Note that Ker Γ is a colocalizing subcategory of T.

Given a subcategory C of a triangulated category T we define full subcategories

⊥C = {X ∈ T | Hom_{T}(X, Σ^{n}Y) = 0 for all Y ∈ C and n ∈ ℤ.}
C^{⊥} = {X ∈ T | Hom_{T}^{∗}(Σ^{n}Y, X) = 0 for all Y ∈ C and n ∈ ℤ}.

Evidently, ^{⊥}C is a localizing subcategory, and C^{⊥} is a colocalizing subcategory. The next lemma summarizes the basic facts about localization and colocalization.

Lemma 2.1. Let T be a triangulated category and S a triangulated subcategory. Then the following are equivalent:

(1) There exists a localization functor L: T → T such that Ker L = S.

(2) There exists a colocalization functor Γ: T → T such that Im Γ = S. In that case both functors are related by a functorial exact triangle

Γ X ⟶ X ⟶ LX ⟶ .

Moreover, there are equalities S^{⊥} = Im L = Ker Γ and ^{⊥}(S^{⊥}) = S.

Proof. See [4, Lemma 3.3]. □

Remark 2.2. There is a dual version of Lemma 2.1 whose formulation is left to the reader. Note that Im F^{op} = Im F and Ker F^{op} = Ker F for any functor F.

Adjoints. We discuss the formal properties of right adjoints of (co)localization functors. This material is the foundation for local homology and cosupport.

Proposition 2.3. Let L, Γ: T → T be exact functors such that L is a localization functor, Γ is a colocalization functor, and both induce a functorial exact triangle Γ X → X → LX →. Then L admits a right adjoint if and only if Γ admits a right adjoint. In that case let Λ and V denote right adjoints of Γ and L, respectively.^{1} Then the following holds.

(1) The functor Λ is a localization functor and V is a colocalization functor. They induce a functorial exact triangle

V X ⟶ X ⟶ ΛX ⟶ .

(2) There are identities

(Im Γ)^{⊥} = Ker Γ = Im L = Im V = Ker Λ = ^{⊥}(Im Λ) .

(3) There are isomorphisms

ΛΓ ⟶^{∼} Λ, Γ ⟶^{∼} Γ Λ, V L ⟶^{∼} L, and V ⟶^{∼} LV .

(4) The functors Γ and Λ induce mutually quasi-inverse equivalences

Im Λ ⟶^{∼} Im Γ and Im Γ ⟶^{∼} Im Λ .

Remark 2.4. The functors L, Γ, Λ, V occurring in the preceding proposition induce the following recollement

\xymatrix{S \ar[r]|{inc} & T \ar[r]|{Q} \ar[l]^{V} \ar[l]_{L} & T/S \ar[l]^{Λ̄} \ar[l]_{Γ̄}}

where S = Im L = Im V and Q: T → T/S denotes the quotient functor so that Γ = Γ̄Q and Λ = Λ̄Q.

Proof of Proposition 2.3. It follows from Proposition A.4 that L admits a right adjoint if and only if there exists a colocalization functor V : T → T with Im V = Im L. Using Lemma 2.1 and the fact that Im L = Ker Γ, it follows that the existence of V is equivalent to the existence of a localization functor Λ: T → T with Ker Λ = Ker Γ. Proposition A.5 and Remark A.6 imply that the existence of Λ is equivalent to the existence of a right adjoint of Γ.

(1) The properties of Λ and V are explained above. The existence of the functorial exact triangle then follows from Lemma 2.1 as Ker Λ = Im V.

1It is customary to write L for a localization functor. A colocalization functor is a localization functor for the opposite category; we denote it Γ, thought of as L turned upside down. The interpretation of local cohomology in the sense of Grothendieck as colocalization provides another reason for the use of Γ. Local homology in the sense of Greenlees and May is denoted Λ; it is a right adjoint of Γ and hence a localization. The corresponding colocalization is thus denoted V.

(2) The identities follow from the first part of the proof and Lemma 2.1.

(3) Combine the localization triangles for L and Λ with the identities in (2).

(4) The isomorphisms in (3) induce isomorphisms

ΛΓ Λ ≅ Λ^{2} ≅ Λ and Γ ΛΓ ≅ Γ^{2} ≅ Γ.

Thus ΛΓ is isomorphic to the identity on Im Λ, while Γ Λ is isomorphic to the identity on Im Γ. □

3. Local cohomology and support

In this section we recall the construction, and basic properties, of local cohomology functors and support for triangulated categories, from [4, 6].

Compact generation. An object C in a triangulated category T admitting setindexed coproducts is compact if the functor Hom_{T}(C, −) commutes with all coproducts. We write T^{c} for the full subcategory of compact objects in T. The category T is compactly generated if it is generated by a set of compact objects.

Recall that we write Σ for the suspension on T. For objects X and Y in T, let

Hom_{T}^{∗}(X, Y) = ⨁_{i∈ℤ} Hom_{T}(X, Σ^{i}Y)

be the graded abelian group of morphisms. Set End_{T}^{∗}(X) = Hom_{T}^{∗}(X, X); this is a graded ring, and Hom_{T}^{∗}(X, Y) is a right End_{T}^{∗}(X) and left End_{T}^{∗}(Y)-bimodule.

Central ring actions. Let R be a graded-commutative ring; thus R is ℤ-graded and satisfies rs = (−1)^{|r||s|}sr for each pair of homogeneous elements r, s in R. We say that a triangulated category T is R-linear, or that R acts on T, if there is a homomorphism ϕ: R → Z^{∗}(T) of graded rings, where Z^{∗}(T) is the graded center of T. This yields for each object X a homomorphism ϕ_{X}: R → End_{T}^{∗}(X) of graded rings such that for all objects X, Y ∈ T the R-module structures on Hom_{T}^{∗}(X, Y) induced by ϕ_{X} and ϕ_{Y} agree, up to the usual sign rule.

Henceforth T will be a compactly generated triangulated category with set-indexed coproducts, and R a graded-commutative noetherian ring acting on T.

Since T is compactly generated with set-indexed coproducts, it follows from the Brown representability theorem that T also admits set-indexed products; see [20, Proposition 8.4.6]. This fact is used without further comment.

Local cohomology and support. We write Spec R for the set of homogeneous prime ideals of R. Fix 𝔭 ∈ Spec R and let M be a graded R-module. The homogeneous localization of M at 𝔭 is denoted by M_{𝔭} and M is called 𝔭-local when the natural map M → M_{𝔭} is bijective.

Given a homogeneous ideal 𝔞 in R, we set

V(𝔞) = {𝔭 ∈ Spec R | 𝔭 ⊇ 𝔞} .

A graded R-module M is 𝔞-torsion if each element of M is annihilated by a power of 𝔞; equivalently, if M_{𝔭} = 0 for all 𝔭 ∈ Spec R ∖ V(𝔞).

The specialization closure of a subset U of Spec R is the set

cl U = {𝔭 ∈ Spec R | there exists 𝔮 ∈ U with 𝔮 ⊆ 𝔭}.

The subset U is specialization closed if cl U = U; equivalently, if U is a union of Zariski closed subsets of Spec R. For each specialization closed subset V of Spec R, we define the full subcategory of T of V-torsion objects as follows:

T_{V} = {X ∈ T | Hom_{T}^{∗}(C, X)_{𝔭} = 0 for all C ∈ T^{c}, 𝔭 ∈ Spec R ∖ V}.

This is a localizing subcategory and there exists a localization functor L_{V}: T → T such that Ker L_{V} = T_{V}; see [4, Lemma 4.3, Proposition 4.5].

The localization functor L_{V} induces a colocalization functor on T, which we denote Γ_{V}, and call the local cohomology functor with respect to V; see Section 2. For each object X in T there is then an exact localization triangle

Γ_{V}X ⟶ X ⟶ L_{V}X ⟶ .

In [4] we established a number of properties of these functors; for instance, that they commute with all coproducts in T, see [4, Corollary 6.5].

For each 𝔭 in Spec R and each object X in T set

X_{𝔭} = L_{Z(𝔭)}X , where Z(𝔭) = {𝔮 ∈ Spec R | 𝔮 ⊈ 𝔭}.

The notation is justified by the fact that, by [4, Theorem 4.7], the adjunction morphism X → X_{𝔭} induces for any compact object C an isomorphism of R-modules

Hom_{T}^{∗}(C, X)_{𝔭} ⟶^{∼} Hom_{T}^{∗}(C, X_{𝔭}) .

We say X is 𝔭-local if the adjunction morphism X → X_{𝔭} is an isomorphism; this is equivalent to the condition that there exists some isomorphism X ≅ X_{𝔭} in T.

Consider the exact functor Γ_{𝔭}: T → T obtained by setting

Γ_{𝔭}X = Γ_{V(𝔭)}(X_{𝔭}) for each object X in T,

and let Γ_{𝔭}T denote its essential image. One has a natural isomorphism Γ_{𝔭}^{2} ≅ Γ_{𝔭}, and an object X from T is in Γ_{𝔭}T if and only if the R-module Hom_{T}^{∗}(C, X) is 𝔭-local and 𝔭-torsion for every compact object C; see [4, Corollary 4.10].

The support of an object X in T is by definition the set

supp_{R} X = {𝔭 ∈ Spec R | Γ_{𝔭}X ≠ 0}.

One has supp_{R} X = ∅ if and only if X = 0 holds; see [4, Theorem 5.2].

Koszul objects. For each object X in T and each homogeneous ideal 𝔞 in R, we denote X//𝔞 a Koszul object on a finite sequence of elements generating the ideal 𝔞; see [4, §5]. Its construction depends on a choice of a generating sequence, but the localizing subcategory generated by it is independent of choice, and depends only on the radical ideal of 𝔞; this follows from [6, Proposition 2.11(2)]. Set

X(𝔭) = (X//𝔭)_{𝔭} for each 𝔭 ∈ Spec R.

The following computations will be used often:

(3.1) supp_{R}(X//𝔭) = V(𝔭) ∩ supp_{R} X and supp_{R} X(𝔭) = {𝔭} ∩ supp_{R} X

For the first one, see [6, Lemma 2.6]; the second follows, given [4, Theorem 5.6].

The first part of the result below is [4, Theorem 6.4], see also [6, Proposition 2.7]; the second one is part of [6, Proposition 3.9].

Theorem 3.2. Suppose G is a set of compact generators for T. For each specialization closed subset V and 𝔭 ∈ Spec R, there are equalities

T_{V} = Loc_{T} (C//𝔭 | C ∈ G and 𝔭 ∈ V) and Γ_{𝔭}T = Loc_{T}(C(𝔭) | C ∈ G),

where both generating sets consist of compact objects. □

4. Local homology and cosupport

Let T denote a compactly generated R-linear triangulated category, as in Section 3. We introduce local homology functors and a notion of cosupport for T.

Local homology. Fix a specialization closed subset V ⊆ Spec R. The functors L_{V} and Γ_{V} on T preserve coproducts by [4, Corollary 6.5] and hence have right adjoints, by Brown representability. Following the notation in Proposition 2.3, this yields adjoint pairs (L_{V}, V^{V}) and (Γ_{V}, Λ^{V}), and, for each X ∈ T, an exact triangle

(4.1)  V^{V}X ⟶ X ⟶ Λ^{V}X ⟶

We call Λ^{V} the local homology functor with respect to V; see Remark 4.17.

The commutation rules for the functors L_{V} and Γ_{V} given in [4, Proposition 6.1] carry over to their right adjoints: For any specialization closed subset W of Spec R there are isomorphisms:

(4.2)  \begin{matrix}Λ^{V}Λ^{W} ≅ Λ^{V∩W} ≅ Λ^{W}Λ^{V} \\ V^{V}V^{W} ≅ V^{V∪W} ≅ V^{W}V^{V} \\ V^{V}Λ^{W} ≅ Λ^{W}V^{V} \end{matrix}

If V ⊇ W holds, then these isomorphisms and Proposition 2.3(2) yield:

(4.3)  V^{V}Λ^{W} ≅ 0 ≅ Λ^{W}V^{V} .

These facts will be used without comment. For each 𝔭 ∈ Spec R set

Λ^{𝔭} = Λ^{V(𝔭)}V^{Z(𝔭)} .

Note that Λ^{𝔭} ≅ V^{Z(𝔭)}Λ^{V(𝔭)}; that Λ^{𝔭} ≅ (Λ^{𝔭})^{2}; and that (Γ_{𝔭}, Λ^{𝔭}) is an adjoint pair.

Cosupport. The cosupport of an object X in T is the set

cosupp_{R} X = {𝔭 ∈ Spec R | Λ^{𝔭}X ≠ 0}.

Cosupport can be computed using Koszul objects, recalled in Section 3.

Proposition 4.4. For each object X in T and 𝔭 ∈ Spec R, one has

𝔭 ∈ cosupp_{R} X ⟺ Hom_{T}(C(𝔭), X) ≠ 0 for some C ∈ T .

Moreover, the object C can be chosen from any set of compact generators for T.

Proof. If Λ^{𝔭}X ≠ 0, then Hom_{T}(Λ^{𝔭}X, Λ^{𝔭}X) ≠ 0, and hence Hom_{T}(Γ_{𝔭}Λ^{𝔭}X, X) ≠ 0, since (Γ_{𝔭}, Λ^{𝔭}) form an adjoint pair. Since Γ_{𝔭}Λ^{𝔭}X is evidently in Γ_{𝔭}T, the last condition implies Hom_{T}(C(𝔭), X) ≠ 0 for some compact object C which is part of a generating set for T, by Theorem 3.2.

Conversely, if Hom_{T}(C(𝔭), X) ≠ 0 for some C ∈ T, then since Γ_{𝔭}C(𝔭) ≅ C(𝔭), by, for instance, Theorem 3.2, one obtains that

Hom_{T}(C(𝔭), Λ^{𝔭}X) ≅ Hom_{T}(Γ_{𝔭}C(𝔭), X) ≅ Hom_{T}(C(𝔭), X) ≠ 0 .

Thus, Λ^{𝔭}X ≠ 0, that is to say, 𝔭 is in cosupp_{R} X. □

A important property of cosupport is that it is non-empty for non-zero objects. We deduce this result from the corresponding statement for supports and the result above. For another perspective, see Theorem 4.13.

Theorem 4.5. For any X ∈ T, one has cosupp_{R} X = ∅ if and only if X = 0.

Proof. Clearly, X = 0 implies cosupp_{R} X = ∅.

If X ≠ 0, then supp_{R} X ≠ ∅, by [4, Theorem 5.2]. Pick a prime 𝔭 in supp_{R} X, maximal with respect to inclusion. Then supp_{R}(X//𝔭) = {𝔭}, by (3.1), which implies, in particular, that X//𝔭 ≠ 0; equivalently, Hom_{T}(X//𝔭, X//𝔭) ≠ 0. Moreover, X//𝔭 is 𝔭-local, hence isomorphic to X(𝔭), which explains the isomorphism below:

Hom_{T}(X(𝔭), X//𝔭) ≅ Hom_{T}(X//𝔭, X//𝔭) ≠ 0 .

Since X//𝔭 is in Thick_{T}(X), by construction, it follows that Hom_{T}(X(𝔭), X) ≠ 0. Hence 𝔭 ∈ cosupp_{R} X, by Proposition 4.4. □

Next we describe further basic properties of local homology functors and cosupports. The one below is immediate from the exactness of the functor Λ^{𝔭}.

Proposition 4.6. For each exact triangle X → Y → Z → ΣX in T, one has

cosupp_{R} Y ⊆ cosupp_{R} X ∪ cosupp_{R} Z and cosupp_{R} ΣX = cosupp_{R} X. □

There is a more precise result for exact triangles (4.1).

Proposition 4.7. Let V be a specialization closed subset of Spec R. For each X in T the following equalities hold:

cosupp_{R} Λ^{V}X = V ∩ cosupp_{R} X
cosupp_{R} V^{V}X = ( Spec R ∖ V) ∩ cosupp_{R} X.

Proof. Fix 𝔭 in Spec R. If 𝔭 ∈ V, then Λ^{V(𝔭)}Λ^{V} = Λ^{V(𝔭)}, by (4.2), and V^{Z(𝔭)}Λ^{V} = 0 if 𝔭 ∉ V, by (4.3). Hence one gets that:

Λ^{𝔭}Λ^{V} = \begin{cases}Λ^{𝔭} & if 𝔭 ∈ V, \\ 0 & otherwise. \end{cases}

The identity for cosupp_{R} Λ^{V}X follows. The proof of the second one is similar. □

The preceding result and Theorem 4.5 yield:

Corollary 4.8. Let V ⊆ Spec R be specialization closed and X an object in T. The following conditions are equivalent:

(1) cosupp_{R} X ⊆ V.

(2) X ∈ Im Λ^{V}.

(3) the natural map X → Λ^{V}X is an isomorphism. □

This result above is complemented by:

Corollary 4.9. Let V ⊆ Spec R be specialization closed and X an object in T. The following conditions are equivalent:

(1) cosupp_{R} X ⊆ Spec R ∖ V.

(1′) supp_{R} X ⊆ Spec R ∖ V.

(2) X ∈ Im V^{V}.

(2′) X ∈ Im L_{V}.

(3) The natural map V^{V}X → X is an isomorphism.

(3′) The natural map X → L_{V}X is an isomorphism. In particular, cosupp_{R} X ⊆ Spec R ∖ Z(𝔭) if and only if X is 𝔭-local.

Proof. The equivalence of (1), (2), and (3) follows from Proposition 4.7 and Theorem 4.5, while the equivalence of (1′), (2′), and (3′) is part of [4, Corollary 5.7]. It remains to note that (2) ⇔ (2′), by Proposition 2.3(2).

The last assertion is (1) ⇔ (3′), applied to V = Z(𝔭). □

The next result is an analogue of [4, Corollary 5.8].

Corollary 4.10. Let X, Y be objects in T. Then cosupp_{R} X ∩ cl(cosupp_{R} Y) = ∅ implies Hom_{T}^{∗}(X, Y) = 0.

Proof. Set V = cl(cosupp_{R} Y). Then X is in Im V^{V}, by Corollary4.9, and Y is in Im Λ^{V}, by Corollary 4.8, so Hom_{T}^{∗}(X, Y) = 0, by Proposition 2.3(2). □

The next goal is Theorem 4.13; the following two results prepare for its proof. The one below is extracted from [4, Lemma 5.11].

Lemma 4.11. Let 𝔞 be a homogeneous ideal in R. For any objects X and Y in T, the following statements hold.

(1) The R-modules Hom_{T}^{∗}(X//𝔞, Y) and Hom_{T}^{∗}(X, Y//𝔞) are 𝔞-torsion.

(2) Hom_{T}^{∗}(X, Y) = 0 implies Hom_{T}^{∗}(X, Y//𝔞) = 0 and the converse holds when

the R-module Hom_{T}^{∗}(X, Y) = 0 is 𝔞-torsion.

(3) Hom_{T}^{∗}(X//𝔞, Y) = 0 if and only if Hom_{T}^{∗}(X, Y//𝔞) = 0. □

The equality below is a version of (3.1) for cosupport.

Lemma 4.12. Let 𝔞 be a homogeneous ideal in R and X an object in T. Then

cosupp_{R}(X//𝔞) = V(𝔞) ∩ cosupp_{R} X .

Proof. From (3.1) one gets an equality

supp_{R}(C(𝔭)//𝔞) = V(𝔞) ∩ {𝔭} ∩ supp_{R} C for any C ∈ T.

If Hom^{∗}(C(𝔭), X//𝔞) ≠ 0 for some C ∈ T, then Hom^{∗}(C(𝔭)//𝔞, X) ≠ 0; this follows from Lemma 4.11(3). In particular, C(𝔭)//𝔞 ≠ 0, so 𝔭 ∈ V(𝔞), by the equality above. One thus obtains from Proposition 4.4 that cosupp_{R}(X//𝔞) ⊆ V(𝔞).

When 𝔭 ∈ V(𝔞) holds, it follows from Lemma 4.11(1), and the observation that C(𝔭) is isomorphic to C_{𝔭}//𝔭, that the R-module Hom_{T}^{∗}(C(𝔭), X) is 𝔭-torsion, and hence also 𝔞-torsion. Therefore, Lemma 4.11(2) yields:

Hom_{T}^{∗}(C(𝔭), X) ≠ 0 ⟺ Hom_{T}^{∗}(C(𝔭), X//𝔞) ≠ 0 .

The desired equality involving cosupport now follows from Proposition 4.4. □

Given U ⊆ Spec R, we write max U for the set of elements 𝔭 ∈ U such that 𝔮 ∈ U and 𝔮 ⊇ 𝔭 imply 𝔮 = 𝔭. Recall that Λ^{𝔭}T denotes the essential image of Λ^{𝔭}.

Theorem 4.13. For each object X in T there is an equality:

max(supp_{R} X) = max(cosupp_{R} X) .

Moreover, cosupp_{R} X(𝔭) ⊆ {𝔭} and X(𝔭) ∈ Λ^{𝔭}T ∩ Γ_{𝔭}T, for each 𝔭 ∈ Spec R.

Proof. We prove max(supp_{R} X) ⊆ cosupp_{R} X and max(cosupp_{R} X) ⊆ supp_{R} X; the first equality would then follow.

Fix 𝔭 in max(supp_{R} X). Then supp_{R}(X//𝔭) = {𝔭}, by (3.1), so X//𝔭 is 𝔭-local; see Corollary 4.9. It is always 𝔭-torsion, so X//𝔭 = X(𝔭), and then

Hom_{T}^{∗}(X(𝔭), X//𝔭) ≅ Hom_{T}^{∗}(X//𝔭, X//𝔭) ≠ 0 .

This implies that 𝔭 is in cosupp_{R}(X//𝔭) by Proposition 4.4, hence also that it is in cosupp_{R} X, by Lemma 4.12.

If 𝔭 ∈ max(cosupp_{R} X), then cosupp_{R}(X//𝔭) = {𝔭}, by Lemma 4.12. Therefore the object X//𝔭 is 𝔭-local, by Corollary 4.9, and 𝔭-torsion, so supp_{R}(X//𝔭) = {𝔭}. It remains to recall (3.1) to conclude that 𝔭 ∈ supp_{R} X.

Finally, cosupp_{R} X(𝔭) ⊆ V(𝔭) by Lemma 4.12, since X(𝔭) ≅ X_{𝔭}//𝔭. Thus the inclusion supp_{R} X(𝔭) ⊆ {𝔭} from (3.1) implies the corresponding inclusion for cosupport. It then follows from Corollaries 4.8 and 4.9 that X(𝔭) is in Λ^{𝔭}T. On the other hand, X(𝔭) is also in Γ_{𝔭}T, by [4, Corollary 5.7]. □

To round off this material, we prove an analogue of Proposition 4.4 for supports.

Proposition 4.14. For each object X in T and 𝔭 ∈ Spec R, one has

𝔭 ∈ supp_{R} X ⟺ Hom_{T}(X, Y (𝔭)) ≠ 0 for some Y ∈ T .

Proof. When 𝔭 is in supp_{R} X it follows from (3.1) that X(𝔭) ≠ 0, hence

Hom_{T}^{∗}(X//𝔭, X(𝔭)) ≅ Hom_{T}^{∗}(X(𝔭), X(𝔭)) ≠ 0 ;

where the first isomorphism holds as X(𝔭) is 𝔭-local. Since X//𝔭 is in Thick_{T}(X), one obtains that Hom_{T}(X, X(𝔭)) ≠ 0. This settles one implication.

The other implication follows from the chain of isomorphisms

Hom_{T}(X, Y (𝔭)) ≅ Hom_{T}(X, Λ^{𝔭}Y (𝔭)) ≅ Hom_{T}(Γ_{𝔭}X, Y (𝔭)),

where the first one holds because Λ^{𝔭}Y (𝔭) ≅ Y (𝔭), by Theorem 4.13. □

Discrete sets. Recall that a subset U ⊆ Spec R is discrete if 𝔭 ⊆ 𝔮 implies 𝔭 = 𝔮 for each pair of primes 𝔭, 𝔮 ∈ U.

Proposition 4.15. Let X and Y be objects of T and U a discrete subset of Spec R. When supp_{R} X ⊆ U and cosupp_{R} Y ⊆ U there are natural isomorphisms

X ⟵^{∼} ∐_{𝔭∈U} Γ_{V(𝔭)}X ⟶^{∼} ∐_{𝔭∈U} Γ_{𝔭}X and Y ⟶^{∼} ∏_{𝔭∈U} Λ^{V(𝔭)}Y ⟵^{∼} ∏_{𝔭∈U} Λ^{𝔭}Y .

Proof. The statement for X is proved in [6, Proposition 3.3]; see also [4, Theorem 7.1]. Modifying the arguments by taking adjunctions, and taking into account Proposition 4.16 below, yields the proof of the statement for Y. □

Analogues of the next statement hold for V^{V} and Λ^{V} also.

Proposition 4.16. Given objects {X_{i}}_{i∈I} in T there is a natural isomorphism

Λ^{𝔭} (∏_{i∈I} X_{i}) ⟶^{∼} ∏_{i∈I} Λ^{𝔭}X_{i}.

In particular, for any subset U ⊆ Spec R the full subcategory with objects

{X ∈ T | cosupp_{R} X ⊆ U}

is a colocalizing subcategory of T.

Proof. Right adjoints distribute over products. This applies to Λ^{𝔭}, which is right adjoint to Γ_{𝔭}, to yield the desired isomorphism. Given this, the second part of the statement follows, for the subcategory in question equals ⋂_{𝔭∉U} Ker Λ^{𝔭}. □

Commutative noetherian rings. Let A be a commutative noetherian ring and D(A) the derived category of the category of all A-modules. The category D(A) is triangulated and compactly generated; indeed, A is a compact generator. It is also A-linear, where for each M ∈ D(A), the homomorphism A → Hom_{D}(M, M) is given by scalar multiplication.

Remark 4.17. Fix an ideal 𝔞 in A. In [4, Theorem 9.1] it is proved that Γ_{V(𝔞)} is the derived functor of the 𝔞-torsion functor, which assigns an A-module M to the module \underrightarrow{lim} Hom_{A}(A/𝔞^{n}, M). Greenlees and May [14, §2], see also Lipman [19, §4], proved that the right adjoint of the latter is local homology and that it coincides with left derived functor of the 𝔞-adic completion functor, which assigns M to \underleftarrow{lim} M/𝔞^{n}M. In commutative algebra literature (as in [19]), the 𝔞-adic completion functor itself would usually be denoted Λ_{𝔞}. Our choice of notation, Λ^{V(𝔞)} for the right adjoint of Γ_{V(𝔞)} is based on this connection.

The functor V^{V} too has, in this context, a familiar avatar, at least in the special case V = Z(𝔭) for some 𝔭 in Spec A: The morphism M → M_{𝔭} = L_{Z(𝔭)}M is the usual localization map, see the proof of [4, Theorem 9.1], so it follows from the classical Hom-tensor adjunction isomorphism that its right adjoint is

V^{Z(𝔭)}M ≅ RHom_{A}(A_{𝔭}, M) ,

and the morphism RHom_{A}(A_{𝔭}, M) → M is the one induced by A → A_{𝔭}. Hence

cosupp_{A} M = {𝔭 ∈ Spec A | RHom_{A}(A_{𝔭}, Λ^{V(𝔭)}M) ≠ 0} .

In practice however, the cosupport seems hard to compute, even for M = A. What little we know about this is contained in the following results.

Proposition 4.18. For each M in D(ℤ) with H^{∗}M finitely generated, one has

supp_{ℤ} M = {𝔭 ∈ Spec ℤ | (H^{∗}M)_{𝔭} ≠ 0} = cosupp_{ℤ} M.

Proof. The equality on the left is easily verified. For the one on the right, given Theorem 4.13, it suffices to prove that the zero ideal is in cosupp_{ℤ} M if and only if it is also in supp_{ℤ} M. By Proposition 4.4, this amounts to verifying that

RHom_{ℤ}(ℚ, M) ≠ 0 ⟺ ℚ ⊗_{ℤ}^{L} M ≠ 0 .

Since M ≅ H^{∗}M in D(ℤ) it suffices to verify the equivalence above when M is an indecomposable finitely generated ℤ-module, hence of the form ℤ/nℤ, for some n ≥ 0. When n ≥ 1, one has RHom_{ℤ}(ℚ, ℤ/nℤ) = 0 = ℚ ⊗_{ℤ}^{L} ℤ/nℤ. It remains to verify that RHom_{ℤ}(ℚ, ℤ) ≠ 0; see [11, §51, Exercise 7]. □

For complete local rings on the other hand, the difference between cosupport and support can be as large as Theorem 4.13 permits.

Proposition 4.19. Let A be a commutative noetherian ring and 𝔞 an ideal in A. The following conditions are equivalent:

(1) A is 𝔞-adically complete.

(2) cosupp_{A} A ⊆ V(𝔞) holds.

(3) cosupp_{A} M ⊆ V(𝔞) holds for any M ∈ D(A) with H^{∗}M finitely generated.

Proof. The equivalence of (1) and (2) is contained in Corollary 4.8. Thus it remains to prove that (1) implies (3).

When A is 𝔞-adically complete so is any finitely generated module. Since the 𝔞-adic completion functor is exact on finitely generated modules, one obtains an isomorphism M ⟶^{∼} Λ^{V(𝔞)}M for any such module M, and hence also for any complex M with H^{∗}M finitely generated. Now apply Corollary 4.8. □

For any object X in T and 𝔭 ∈ Spec R, the support of X_{𝔭} is contained in that of X; see [4, Theorem 5.6]. The corresponding statement for cosupports does not hold, which speaks to one of the difficulties in computing this invariant.

Example 4.20. Let (A, 𝔪) be a local ring that is 𝔪-adically complete. Then for any 𝔭 ∈ Spec A ∖ {𝔪} one has

{𝔭} ⊆ cosupp_{A}(A_{𝔭}) ⊈ cosupp_{A} A = {𝔪} .

Indeed, the equality is by Proposition 4.19. That 𝔭 is in cosupp_{A}(A_{𝔭}) can be checked directly, or via Theorem 4.13 and the equality supp_{A}(A_{𝔭}) = Spec A_{𝔭}.

Remark 4.21. The notion of cosupport for an A-module M is not the same as the one introduced by Richardson [23], which is denoted coSupp M. For instance, it follows from [23, Theorem 2.7(vi)] that coSupp ℤ = Spec ℤ ∖ {(0)}, while Proposition 4.18 yields cosupp_{ℤ} ℤ = Spec ℤ. When (A, 𝔪) is local, each non-zero finitely generated module M satisfies coSupp M = {𝔪}, by [23, Theorem 2.7(i)], while Proposition 4.19 implies cosupp_{A} M = {𝔪} if and only if M is 𝔪-adically complete.

5. 𝔭-local and 𝔭-complete objects

As before, let T denote a compactly generated R-linear triangulated category. In this section we investigate the subcategories Λ^{𝔭}T, for each 𝔭 ∈ Spec R. They contain local information about T and are important in classifying its colocalizing subcategories; see the discussion in the last part of this section.

We begin by noting that an object X is in Λ^{𝔭}T if and only if cosupp_{R} X ⊆ {𝔭}, if and only if X is 𝔭-local (that is, X → X_{𝔭} is an isomorphism) and also 𝔭-complete, meaning that the natural map X → Λ^{V(𝔭)}X is an isomorphism; see Corollaries 4.9 and 4.8. Also Λ^{𝔭}T is a colocalizing subcategory of T, by Proposition 4.16.

Recall that X is in Γ_{𝔭}T if and only if supp_{R} X ⊆ {𝔭}, if and only if X is 𝔭-local and 𝔭-torsion; see [4, Corollaries 4.9, 5.10], and that Γ_{𝔭}T is a localizing subcategory.

Dwyer Greenlees correspondence. Fix a prime 𝔭 in Spec R. The result below, which establishes an equivalence between the category of 𝔭-local and 𝔭-complete objects and the category of 𝔭-local and 𝔭-torsion objects, may be viewed as extension of such an equivalence discovered by Dwyer and Greenlees [9, Theorem 2.1], see also Hovey, Palmieri, and Strickland [15, Theorem 3.3.5], to our setting.

Proposition 5.1. The functors Γ_{𝔭}: Λ^{𝔭}T → Γ_{𝔭}T and Λ^{𝔭}: Γ_{𝔭}T → Λ^{𝔭}T form an adjoint pair, that are mutually quasi-inverse to each other.

Proof. Apply the identities in Proposition 2.3. □

As a triangulated category, Γ_{𝔭}T is compactly generated by {C(𝔭) | C ∈ T^{c}}; see Theorem 3.2. Given Theorem 4.13 and the equivalence Γ_{𝔭}T ⟶^{∼} Λ^{𝔭}T in Proposition 5.1, it follows that the same set generates Λ^{𝔭}T, where the coproduct in Λ^{𝔭}T is the one induced from Γ_{𝔭}T.

Next we describe a set of cogenerators for Λ^{𝔭}T.

Perfect cogeneration. Let U be a triangulated category with set-indexed products. A set of objects S perfectly cogenerates U if the following conditions hold:

(1) If X is an object in U and Hom_{U}(X, S) = 0 for all S ∈ S then X = 0.

(2) If a countable family of maps X_{i} → Y_{i} in U is such that

Hom_{U}(Y_{i}, S) → Hom_{U}(X_{i}, S)

is surjective for all i and all S ∈ S, then so are the induced maps:

Hom_{U}(∏_{i} Y_{i}, S) → Hom_{U}(∏_{i} X_{i}, S) .

Proposition 5.2. If a set of objects S perfectly cogenerates U then Coloc_{U}(S) = U.

Proof. The proof is akin to that of the corollary in [17, §1]. □

Injective objects. Let now T be an R-linear triangulated category as before. For each compact object C in T and each injective R-module I, Brown representability yields an object T_{C}(I) in T and a natural isomorphism:

(5.3)  Hom_{T}(−, T_{C}(I)) ≅ Hom_{R}(Hom_{T}^{∗}(C, −), I).

For 𝔭 ∈ Spec R, let I(𝔭) denote the injective envelope of R/𝔭.

Proposition 5.4. Fix 𝔭 ∈ Spec R. For each compact object C in T, one has

cosupp_{R} T_{C}(I(𝔭)) = supp_{R} C ∩ {𝔮 ∈ Spec R | 𝔮 ⊆ 𝔭} .

Moreover, the set {T_{C//𝔭}(I(𝔭)) | C ∈ T^{c}} perfectly cogenerates Λ^{𝔭}T.

Proof. The shifts of I(𝔭) form a set of injective cogenerators for the category of 𝔭-local R-modules. For any 𝔮 ∈ Spec R and object D in T, one has equivalences

Hom_{T}^{∗}(D(𝔮), T_{C}(I(𝔭))) ≠ 0 ⟺ Hom_{R}^{∗}(Hom_{T}^{∗}(C, D(𝔮)), I(𝔭)) ≠ 0
⟺ Hom_{T}^{∗}(C, D(𝔮)) ≠ 0 and 𝔮 ⊆ 𝔭 .

The first one is by (5.3); the second holds as the R-module Hom_{T}^{∗}(C, D(𝔮)) is 𝔮-local and 𝔮-torsion, by (3.1). Propositions 4.4 and 4.14 now yield the stated equality.

Let X be a non-zero object in Λ^{𝔭}T, so that cosupp_{R} X = {𝔭}, and pick a compact object C with Hom_{T}^{∗}(C(𝔭), X) ≠ 0; see Proposition 4.4. It then follows that:

Hom_{T}^{∗}(C//𝔭, X) ≅ Hom_{T}^{∗}(C//𝔭, Λ^{𝔭}X)

≅ Hom_{T}^{∗}(Γ_{𝔭}(C//𝔭), X)

≅ Hom_{T}^{∗}(C(𝔭), X)

≠ 0

Replacing C by an appropriate suspension Σ^{n}C, if necessary, from the computation above and (5.3) one gets

Hom_{T}(X, T_{C//𝔭}(I(𝔭))) ≅ Hom_{R}(Hom_{T}^{∗}(C//𝔭, X), I(𝔭)) ≠ 0 .

The other condition for perfect cogeneration holds because, for any map X → Y in Λ^{𝔭}T, the induced map

Hom_{T}^{∗}(Y, T_{C//𝔭}(I(𝔭))) → Hom_{T}^{∗}(X, T_{C//𝔭}(I(𝔭)))

is surjective if and only if Hom_{T}^{∗}(C//𝔭, X) → Hom_{T}^{∗}(C//𝔭, Y) is injective. □

Classifying colocalizing subcategories. We say that the local-global principle holds for colocalizing subcategories of T if for each object X in T there is an equality

Coloc_{T}(X) = Coloc_{T}({Λ^{𝔭}X | 𝔭 ∈ Spec R}).

The corresponding notion for localizing subcategories is investigated in [6, §3]. The reformulation below of the local-global principle is easy to prove.

Lemma 5.5. The local-global principle for colocalizing subcategories is equivalent to the statement: For any X ∈ T and any colocalizing subcategory S of T, one has

X ∈ S ⟺ Λ^{𝔭}X ∈ S for each 𝔭 ∈ Spec R.  □

Colocalizing subcategories of T are related to subsets of Spec R via maps

\{ \begin{matrix}colocalizing \\ subcategories of T \end{matrix} \} [figure: σ; τ]\{ \begin{matrix}families (S(𝔭))_{𝔭∈Spec R} with S(𝔭) \\ a colocalizing subcategory of Λ^{𝔭}T \end{matrix} \}

which are defined by σ(S) = (S ∩ Λ^{𝔭}T) and τ(S(𝔭)) = Coloc_{T} (S(𝔭) | 𝔭 ∈ Spec R).

Proposition 5.6. If the local-global principle for colocalizing subcategories of T holds, then the map σ is bijective, with inverse τ.

Proof. We use the fact that Λ^{𝔭} is an idempotent exact functor preserving products. First observe that for each colocalizing subcategory S of T there is an inclusion

(∗)  S ∩ Λ^{𝔭}T ⊆ Λ^{𝔭}S for each 𝔭 ∈ Spec R.

We prove that στ is the identity, that is to say, that for any family (S(𝔭))_{𝔭∈Spec R} of colocalizing subcategories with S(𝔭) ⊆ Λ^{𝔭}T the colocalizing subcategory cogenerated by all the S(𝔭), call it S, satisfies

S ∩ Λ^{𝔭}T = S(𝔭) for each 𝔭 ∈ Spec R.

Note that Λ^{𝔭}S = S(𝔭) holds, since Λ^{𝔭}Λ^{𝔮} = 0 when 𝔭 ≠ 𝔮. Therefore (∗) yields an inclusion S ∩ Λ^{𝔭}T ⊆ S(𝔭). The reverse inclusion is obvious.

For any localizing subcategory S of T, the reformulation of the local-global principle in Lemma 5.5 gives S = Coloc_{T}(S ∩ Λ^{𝔭}T | 𝔭 ∈ Spec R). Thus τσ = id. □

Remark 5.7. In analogy with the notion of stratification for localizing subcategories of T introduced in [6, §4], we say that T is costratified by R if

(C1) The local-global principle holds for colocalizing subcategories of T; (C2) For each 𝔭 ∈ Spec R, the colocalizing subcategory Λ^{𝔭}T contains no proper

non-zero colocalizing subcategories. It is immediate from Proposition 5.6 that when these conditions hold, the maps σ and τ induce a bijection

\{ \begin{matrix}Colocalizing \\ subcategories of T \end{matrix} \} [figure: ^{cosupp^{R}}; _{cosupp{−1 \atop R}}]\{ subsets of supp_{R} T \} .

For the main results of this work, it suffices to consider a version of costratification for tensor triangulated categories, see Section 9; so we do not study the general notion in any great detail.

6. Axioms for support and cosupport

In this section we give an axiomatic description of cosupport, analogous to the one for support in [4, Theorem 5.15]; see also Theorem 6.4 below. This material is not used elsewhere in this paper.

As before, we fix a compactly generated R-linear triangulated category T. The starting point is the following cohomological addendum to Corollary 4.9.

Lemma 6.1. Let V ⊆ Spec R be a specialization closed subset. For each object X in T, the following conditions are equivalent:

(1) V ∩ cosupp_{R} X = ∅.

(1′) V ∩ supp_{R} X = ∅.

(2) Hom_{T}^{∗}(C//𝔭, X) = 0 for all C ∈ T^{c} and 𝔭 ∈ V.

(3) Hom_{T}^{∗}(C, X) is either zero or not 𝔭-torsion for each C ∈ T^{c} and 𝔭 ∈ V.

Proof. (1) ⇔ (1′) is part of Corollary 4.9, while (1′) ⇔ (2) is a consequence of Theorem 3.2. To prove that (2) ⇔ (3), use Lemma 4.11. □

Remark 6.2. As noted above, for each specialization closed subset V of Spec R there is an equality of subcategories

{X ∈ T | cosupp_{R} X ∩ V = ∅} = {X ∈ T | supp_{R} X ∩ V = ∅} .

The subcategory on the left is colocalizing, by Proposition 4.16, while the one on the right is localizing, since Γ_{𝔭} preserves set-indexed coproducts; see [4, Corollary 6.6].

Given a specialization closed subset V ⊆ Spec R, Lemma 6.1 yields cohomological criteria for the (co)support of any object X to be contained in Spec R ∖ V. This leads to axiomatic descriptions of cosupport and support.

Theorem 6.3. There exists a unique assignment sending each object X in T to a subset cosupp_{R} X of Spec R such that the following properties hold:

(1) Cohomology: For each object X in T and each 𝔭 in Spec R, one has

V(𝔭) ∩ cosupp_{R} X ≠ ∅

if and only if Hom_{T}^{∗}(C, X) is non-zero and 𝔭-torsion for some C in T^{c}.

(2) Orthogonality: For objects X and Y in T, one has that

cosupp_{R} X ∩ cl(cosupp_{R} Y) = ∅ implies Hom_{T}(X, Y) = 0.

(3) Exactness: For every exact triangle X → Y → Z → in T, one has

cosupp_{R} Y ⊆ cosupp_{R} X ∪ cosupp_{R} Z.

(4) Separation: For any specialization closed subset V of Spec R and object X in T, there exists an exact triangle X′ → X → X′′ → in T such that

cosupp_{R} X′ ⊆ Spec R ∖ V and cosupp_{R} X′′ ⊆ V.

Proof. Lemma 6.1 implies (1). Corollary 4.10 is (2), and Proposition 4.6 is (3). Proposition 4.7 implies (4). Here one uses for any specialization closed subset V of Spec R the localization triangle (4.1).

Now let σ: T → Spec R be a map satisfying properties (1)–(4).

Fix a specialization closed subset V ⊆ Spec R and an object X ∈ T. It suffices to verify that the following equalities hold:

(∗)  σ(Λ^{V}X) = σ(X) ∩ V and σ(V^{V}X) = σ(X) ∩ (Spec R ∖ V).

Indeed, for any point 𝔭 in Spec R one then obtains that

σ(Λ^{𝔭}X) = σ(V^{Z(𝔭)}Λ^{V(𝔭)}X)
= σ(Λ^{V(𝔭)}X) ∩ (Spec R ∖ Z(𝔭))
= σ(X) ∩ V(𝔭) ∩ (Spec R ∖ Z(𝔭))
= σ(X) ∩ {𝔭}.

Therefore, 𝔭 ∈ σ(X) if and only if σ(Λ^{𝔭}X) ≠ ∅; this last condition is equivalent to Λ^{𝔭}X ≠ 0, by the cohomology property. The upshot is that 𝔭 ∈ σ(X) if and only if 𝔭 ∈ supp_{R} X, which is the desired conclusion.

It thus remains to prove (∗).

Let X′ → X → X′′ → be the triangle associated to V, provided by property (4). It suffices to verify the following statements:

(i) σ(X′′) = σ(X) ∩ V and σ(X′) = σ(X) ∩ (Spec R ∖ V); (ii) Λ^{V}X ≅ X′ and L_{V}X ≅ X′′. The equalities in (i) are immediate from properties (3) and (4). In verifying (ii),

the crucial observation is that, by the cohomology property, for any Y in T one has

V ∩ σ(Y) = ∅ ⟺ Y ∈ Im V^{V} .

Thus X′ is in Im V^{V}. On the other hand, property (2) and Lemma 2.1 imply that X′′ is in Im Λ^{V}. One thus obtains the following morphism of triangles

[figure: X′; α; X; X′′; β; V^{V}X; X; Λ^{V}X]

where the object Cone(α) ≅ Cone(Σ^{−1}β) belongs to Im V^{V} ∩Im Λ^{V}, hence is trivial. Therefore, α and β are isomorphisms, which yields (ii). □

The following axiomatic description of support is a slight modification of [4, Theorem 5.15]. Note that conditions (1) and (3) coincide for support and cosupport. The crucial difference appears in conditions (2) and (4).

Theorem 6.4. There exists a unique assignment sending each object X in T to a subset supp_{R} X of Spec R such that the following properties hold:

(1) Cohomology: For each object X in T and each 𝔭 in Spec R, one has

V(𝔭) ∩ supp_{R} X ≠ ∅

if and only if Hom_{T}^{∗}(C, X) is non-zero and 𝔭-torsion for some C in T^{c}.

(2) Orthogonality: For objects X and Y in T, one has that

cl(supp_{R} X) ∩ supp_{R} Y = ∅ implies Hom_{T}(X, Y) = 0.

(3) Exactness: For every exact triangle X → Y → Z → in T, one has

supp_{R} Y ⊆ supp_{R} X ∪ supp_{R} Z.

(4) Separation: For any specialization closed subset V of Spec R and object X in T, there exists an exact triangle X′ → X → X′′ → in T such that

supp_{R} X′ ⊆ V and supp_{R} X′′ ⊆ Spec R ∖ V.

Proof. Adapt the proof of [4, Theorem 5.15], using Lemma 6.1. □

7. Change of rings and categories

In this section we discuss how support and cosupport is affected by the change of rings and categories. Throughout R is a graded-commutative noetherian ring.

Linear functors. Let T and U be R-linear triangulated categories. We say that a functor F: T → U is R-linear if it is an exact functor such that for each X in T the following diagram is commutative:

\xymatrix{ & R \ar[dr]^{ϕ_{F X}} \ar[dl]_{ϕ_{X}} \\ End_{T}^{∗}(X) \ar[rr]^{F} &  & End_{U}^{∗}(F X)}

Let F: T → U be an R-linear functor, and let X and Y be objects in T and U, respectively. The structure homomorphisms ϕ_{X} and ϕ_{F X} provide two R-module structures on Hom_{U}^{∗}(F X, Y), and the R-linearity of F is equivalent to the claim that these coincide.

We are grateful to the referee for suggesting the following lemma.

Lemma 7.1. Let F: T → U be an R-linear functor and G a right adjoint. Then the following statements hold.

(1) The adjunction isomorphism Hom_{T}^{∗}(X, GY) ⟶^{∼} Hom_{U}^{∗}(F X, GY) is R-linear.

(2) The functor G is R-linear.

Proof. For (1), observe that the adjunction isomorphism can be factored as

Hom_{T}^{∗}(X, GY) ⟶^{F} Hom_{U}^{∗}(F X, F GY) ⟶^{(F X,θY)} Hom_{U}^{∗}(F X, Y)

where the second map is induced by the counit θ: F G → Id_{U}. For (2), note that the map Hom_{U}(Y, Y) ⟶^{G} Hom_{T}(GY, GY) can be factored as

Hom_{U}^{∗}(Y, Y) ⟶^{(θY,Y)} Hom_{U}^{∗}(F GY, Y) ⟶^{∼} Hom_{T}^{∗}(GY, GY)

where the second map is the adjunction isomorphism. □

Proposition 7.2. Let F: T → U be a functor between compactly generated R-linear triangulated categories which preserves set-indexed coproducts and products. Let E be a left adjoint of F, and suppose that F or E is R-linear. For any specialization closed subset V of Spec R there are then natural isomorphisms

F Γ_{V} ≅ Γ_{V}F, F L_{V} ≅ L_{V}F, Γ_{V}E ≅ EΓ_{V}, and L_{V}E ≅ EL_{V} .

Note that the functor F admits a left adjoint by Brown representability, because F preserves set-indexed products.

Proof. For each object X in T, one has an exact triangle

F Γ_{V}X ⟶ F X ⟶ F L_{V}X ⟶

induced by the localization triangle for V. It thus suffices to verify that F Γ_{V}X is in Im Γ_{V} and that F L_{V}X is in Im L_{V}; see [4, §4]. We use the adjunction isomorphisms

Hom_{U}^{∗}(C, F Γ_{V}X) ≅ Hom_{T}^{∗}(EC, Γ_{V}X)
Hom_{U}^{∗}(C, F L_{V}X) ≅ Hom_{T}^{∗}(EC, L_{V}X)

(7.3)

which are R-linear because F or E is R-linear; see Lemma 7.1. Note that EC is compact if C is compact, since F preserves set-indexed coproducts.

An object Y in U belongs to Im Γ_{V} if and only if Hom_{U}^{∗}(C, Y)_{𝔭} = 0 for all compact C ∈ U and all 𝔭 ∈ Spec R ∖ V. Applying this characterization to F Γ_{V}X and Γ_{V}X, the adjunction (7.3) implies that F Γ_{V}X is in Im Γ_{V}.

An object Y in U belongs to Im L_{V} if and only if Hom_{U}^{∗}(C//𝔭, Y) = 0 for all compact C ∈ U and all 𝔭 ∈ V, by Corollary 4.9 and Lemma 6.1. Applying this characterization to F L_{V}X and L_{V}X, the adjunction (7.3) implies that F L_{V}X is in Im L_{V}. Here, one uses that E(C//𝔭) ≅ (EC)//𝔭, and this completes the proof of the first pair of isomorphisms.

The isomorphism F L_{Z(𝔭)} ≅ L_{Z(𝔭)}F implies F (X(𝔭)) ≅ (F X)(𝔭) for each X in T and each 𝔭 in Spec R. Given this, the proof of the isomorphisms involving E is similar: For each object Y in U, it follows from Proposition 4.14 and adjunction that there are inclusions

supp_{R} EΓ_{V}Y ⊆ V and supp_{R} EL_{V}Y ⊆ Spec R ∖ V .

Thus EΓ_{V}Y is in Im Γ_{V} and EL_{V}Y in Im L_{V}. □

Remark 7.4. In the preceding proof, the assumption on F or E to be R-linear is only used for the R-linearity of the adjunction isomorphisms (7.3).

Change of rings. Let S be a graded-commutative noetherian ring, and U an Slinear triangulated category. Given a homomorphism of rings α: R → S, there is a natural R-linear structure on U induced by homomorphisms

R ⟶^{α} S ⟶^{ϕ_{X}} End_{U}^{∗}(X) for X ∈ U.

As usual, α induces a map α^{∗}: Spec S → Spec R, with α^{∗}(𝔮) = α^{−1}(𝔮) for each 𝔮 in Spec S. Observe that if V ⊆ Spec R is specialization closed, then so is the subset (α^{∗})^{−1}V of Spec S.

Proposition 7.5. Let α: R → S be a homomorphism of rings, and U an S-linear triangulated category, with induced R-linear structure via α. Let V ⊆ Spec R be a specialization closed set and W = (α^{∗})^{−1}V. Then there are isomorphisms

Γ_{V} ≅ Γ_{W} and L_{V} ≅ L_{W}.

Proof. It suffices to prove that any object X ∈ U is in U_{V} if and only if it is in U_{W}. Thus one needs to show for any compact object C ∈ U that Hom_{U}^{∗}(C, X)_{𝔭} = 0 for all 𝔭 ∈ Spec R ∖ V if and only if Hom_{U}^{∗}(C, X)_{𝔮} = 0 for all 𝔮 ∈ Spec S ∖ W. This one finds in Lemma 7.6 below. □

Lemma 7.6. Let α: R → S be a homomorphism of graded-commutative noetherian rings and V ⊆ Spec R a specialization closed subset. Given an S-module M, one has M_{𝔭} = 0 for all 𝔭 ∈ Spec R ∖ V if and only if M_{𝔮} = 0 for all 𝔮 ∈ Spec S ∖ (α^{∗})^{−1}V.

Proof. Suppose first that M_{𝔭} = 0 for all 𝔭 ∉ V, and choose 𝔮 ∉ (α^{∗})^{−1}V. Then M_{𝔮} ≅ (M_{α^{∗}(𝔮)})_{𝔮} = 0.

Assume now that M_{𝔭} ≠ 0 for some 𝔭 ∉ V. We view M_{𝔭} as an S_{𝔭}-module and find therefore a prime ideal 𝔮 in Spec S_{𝔭} ⊆ Spec S such that (M_{𝔭})_{𝔮} ≅ M_{𝔮} is non-zero. It remains to observe that α^{∗}(𝔮) ⊆ 𝔭 and hence that α^{∗}(𝔮) is not in V. □

Change of rings and categories. Henceforth, we say (F; α): (T; R) ⟶ (U; S) is an exact functor to mean that T and U are compactly generated R-linear and S-linear triangulated categories, respectively; α: R → S is a homomorphism of graded rings; and F is an exact functor that is R-linear with respect to the induced R-linear structure on U; in other words, that the diagram

\xymatrix{R \ar[r]^{α} \ar[d]_{ϕ_{X}} & S \ar[d]^{ϕ_{F X}} \\ End_{T}^{∗}(X) \ar[r]^{F} & End_{U}^{∗}(F X)}

is commutative for each X ∈ T.

Theorem 7.7. Let (F; α): (T; R) → (U; S) be an exact functor which preserves set-indexed coproducts and products. Let E be a left adjoint and G a right adjoint of F. Let V ⊆ Spec R be a specialization closed subset and set W = (α^{∗})^{−1}V. Then there are natural isomorphisms:

F Γ_{V} ≅ Γ_{W}F, F L_{V} ≅ L_{W}F, Γ_{V}E ≅ EΓ_{W}, and L_{V}E ≅ EL_{W} (1)

F Λ^{V} ≅ Λ^{W}F, F V^{V} ≅ V^{W}F, Λ^{V}G ≅ GΛ^{W}, and V^{V}G ≅ GV^{W} . (2)

This result contains Propositions 7.2 and 7.5: to recover the first, set α = id_{R}; for the second set F = Id_{T}. On the other hand, it is proved using the latter results.

Proof. Since F is linear with respect to the induced R-linear structure on U, Propositions 7.2 and 7.5 yield the following isomorphism:

F Γ_{V} ≅ Γ_{V}F ≅ Γ_{W}F.

The other isomorphisms in (1) can be obtained in the same way.

The isomorphisms in (2) are obtained by taking right adjoints of those in (1). □

As applications, we establish results which track the change in support along linear functors; this is one reason we have had to introduce these notions.

Corollary 7.8. Let (F; α): (T; R) → (U; S) be an exact functor which preserves set-indexed coproducts and products. Let E be a left adjoint and G a right adjoint of F. Then for X ∈ T and Y ∈ U there are inclusions:

(1) α^{∗}(supp_{S} F X) ⊆ supp_{R} X and supp_{R} EY ⊆ α^{∗}(supp_{S} Y)

(2) α^{∗}(cosupp_{S} F X) ⊆ cosupp_{R} X and cosupp_{R} GY ⊆ α^{∗}(cosupp_{S} Y) .

Each inclusion is an equality when the corresponding functor is faithful on objects.

Proof. Let 𝔭 be a point in Spec R, and pick specialization closed subsets V and W of Spec R such that {𝔭} = V ∖ W. For example, set V = V(𝔭) and W = V ∖ {𝔭}.

Setting Ṽ = (α^{∗})^{−1}V and W̃ = (α^{∗})^{−1}W, one gets isomorphisms

F (Γ_{𝔭}X) = F L_{W}Γ_{V}X = LW̃ΓṼF X .

Observing that Ṽ ∖ W̃ = (α^{∗})^{−1}{𝔭}, this yields equalities

supp_{S} F (Γ_{𝔭}X) = supp_{S} F X ∩ (Ṽ ∖ W̃) = supp_{S} F X ∩ (α^{∗})^{−1}{𝔭} .

Thus, α^{∗}(supp_{S} F X) ⊆ supp_{R} X, and equality holds if F is faithful on objects. The other inclusions are obtained in the same way. □

In the preceding result, the stronger conclusion supp_{S} F X = (α^{∗})^{−1} supp_{R} X need not hold, even when F is an equivalence of categories.

Example 7.9. Let R be a field, set S = R[a]/(a^{2} − a), and let U = D(S) denote the derived category of S-modules, with canonical S-linear structure. Let T = U and view this as an R-linear triangulated category via the inclusion α: R → S.

Let F: T → U be the identity functor; it is evidently compatible with α and faithful. Observe however that for the module X = S/(a) in T one has

supp_{R} X = Spec R and Spec_{S} F X = {(a)}.

On the other hand, (α^{∗})^{−1} supp_{R} X = Spec S.

Corollary 7.10. Let (F; α): (T; R) → (U; S) be an exact functor which preserves set-indexed coproducts and products. Let E be a left adjoint and G a right adjoint of F. Let 𝔭 ∈ Spec R and suppose that U = (α^{∗})^{−1}{𝔭} is a discrete subset of Spec S. Then there are isomorphisms:

F Γ_{𝔭} ≅ ∐_{𝔮∈U} Γ_{𝔮}F and Γ_{𝔭}E ≅ ∐_{𝔮∈U} EΓ_{𝔮} (1)

F Λ^{𝔭} ≅ ∏_{𝔮∈U} Λ^{𝔮}F and Λ^{𝔭}G ≅ ∏_{𝔮∈U} GΛ^{𝔮} . (2)

Proof. Choose specialization closed subsets V and W of Spec R with V ∖ W = {𝔭}. Setting Ṽ = (α^{∗})^{−1}V and W̃ = (α^{∗})^{−1}W, one gets isomorphisms

F Γ_{𝔭} = F L_{W}Γ_{V} ≅ LW̃ΓṼF ≅ ∐_{𝔮∈U} Γ_{𝔮}L_{W}Γ_{V}F ≅ ∐_{𝔮∈U} Γ_{𝔮}F ,

where the first one follows Theorem 7.7. The second is by Proposition 4.15, which applies since Ṽ ∖ W̃ = U and U is discrete. The last isomorphism holds because one has Γ_{𝔮}LW̃ΓṼ ≅ Γ_{𝔮} for all 𝔮 ∈ U; see [4, Proposition 6.1].

The other isomorphisms can be obtained in the same way. □

The next result involves a notion of costratification of R-linear triangulated categories. This has been introduced in Remark 5.7, and the analogous notion of stratification is from [6, §4].

Theorem 7.11. Let (F; α): (T; R) → (U; S) be an exact functor which preserves set-indexed coproducts and products, and fix an object X in T.

(1) If T is costratified by R and the right adjoint of F is faithful on objects, then

supp_{S} F X = (α^{∗})^{−1}(supp_{R} X) ∩ supp_{S} U.

(2) If T is stratified by R and the left adjoint of F is faithful on objects, then

cosupp_{S} F X = (α^{∗})^{−1}(cosupp_{R} X) ∩ supp_{S} U.

Proof. We prove the statement concerning cosupports; the argument for the one for supports is exactly analogous.

To begin with, from Corollary 7.8 one gets an inclusion

cosupp_{S} F X ⊆ (α^{∗})^{−1}(cosupp_{R} X) ∩ cosupp_{S} U

Now fix a 𝔮 ∈ supp_{S} U with 𝔮 ∉ cosupp_{S} F X. We need to show that 𝔭 = α^{∗}(𝔮) is not in cosupp_{R} X. Let E be a left adjoint of F. Using adjunction, one has

Hom_{T}(EΓ_{𝔮}−, X) ≅ Hom_{U}(−, Λ^{𝔮}F X) = 0.

There exists some object U in U such that EΓ_{𝔮}U ≠ 0, since 𝔮 ∈ supp_{S} U and E is faithful on objects. Moreover, EΓ_{𝔮}U belongs to Γ_{𝔭}T, by Corollary 7.8. Since R stratifies T, the subcategory Γ_{𝔭}T contains no non-trivial localizing subcategories, and hence coincides with Loc_{T}(EΓ_{q}U). Thus

0 = Hom_{T}(Γ_{𝔭}−, X) ≅ Hom_{T}(−, Λ^{𝔭}X),

and therefore 𝔭 ∉ cosupp_{R} X. □

Perfect generators and cogenerators. For any subset U of Spec R we consider the full subcategories

T_{U} = {X ∈ T | supp_{R} X ⊆ U} and T^{U} = {X ∈ T | cosupp_{R} X ⊆ U} .

The notion of a set of perfect cogenerators was recalled in Section 5. The notion of a set of perfect generators is analogous; see [17].

Lemma 7.12. Let (F; α): (T; R) → (U; S) be an exact functor which preserves setindexed coproducts and products. Let U be a subset of Spec R and set Ũ = (α^{∗})^{−1}U. Then the following statements hold:

(1) When F is faithful on objects from T_{U}, its left adjoint maps any set of perfect generators of UŨ to a set of perfect generators of T_{U}.

(2) When F is faithful on objects from T^{U}, its right adjoint maps any set of perfect cogenerators of UŨ to a set of perfect cogenerators of T^{U},

Proof. Let E denote a left adjoint of F. It follows from Corollary 7.8 that F and E restrict to functors between T_{U} and UŨ. Now use adjunction to prove (1). The proof of (2) is analogous. □

8. Tensor triangulated categories

In this section we discuss special properties of triangulated categories which hold when they have a tensor structure.

Let (T, ⊗, 𝟙) be a tensor triangulated category as defined in [4, §8]. In particular, T is a compactly generated triangulated category with a symmetric monoidal structure; ⊗ is its tensor product and 𝟙 the unit of the tensor product. The tensor product is exact in each variable and preserves coproducts.

By Brown representability there are function objects Hom(X, Y) satisfying

(8.1)  Hom_{T}(X ⊗ Z, Y) ≅ Hom_{T}(Z, Hom(X, Y)),

and we write X^{∨} for the Spanier–Whitehead dual Hom(X, 𝟙). Note that the adjunction extends to function objects, in the sense that there are natural isomorphisms

Hom(X ⊗ Z, Y) ≅ Hom(Z, Hom(X, Y)) .

This is an easy consequence of Yoneda’s lemma.

We shall assume that the tensor unit 𝟙 is compact and that all compact objects C are strongly dualizable in the sense that the canonical morphism

C^{∨} ⊗ X → Hom(C, X)

is an isomorphism for all X in T. We also assume that Hom(−, Y) is exact for each object Y in T.^{2}

Canonical actions. The symmetric monoidal structure of T ensures that the endomorphism ring End_{T}^{∗}(𝟙) is graded commutative. It acts on T via homomorphisms

End_{T}^{∗}(𝟙) ⟶^{X⊗−} End_{T}^{∗}(X).

In particular, any homomorphism R → End_{T}^{∗}(𝟙) of rings with R graded commutative induces an action of R on T. We say that an R action on T is canonical if it arises from such a homomorphism. In that case there are for each specialization closed subset V and point 𝔭 of Spec R natural isomorphisms

(8.2)  Γ_{V}X ≅ X ⊗ Γ_{V}𝟙, L_{V}X ≅ X ⊗ L_{V}𝟙, and Γ_{𝔭}X ≅ X ⊗ Γ_{𝔭}𝟙.

These isomorphisms are from [4, Theorem 8.2, Corollary 8.3].^{3}

Proposition 8.3. Let V ⊆ Spec R be a specialization closed subset and 𝔭 ∈ Spec R. Given objects X and Y in T, there are natural isomorphisms

Hom(Γ_{V}X, Y) ≅ Hom(X, Λ^{V}Y) ≅ Λ^{V} Hom(X.Y),
Hom(L_{V}X, Y) ≅ Hom(X, V^{V}Y) ≅ V^{V} Hom(X, Y),
Hom(Γ_{𝔭}X, Y) ≅ Hom(X, Λ^{𝔭}Y) ≅ Λ^{𝔭} Hom(X, Y).

In particular, there are natural isomorphisms

Λ^{V}X ≅ Hom(Γ_{V}𝟙, X), V^{V}X ≅ Hom(L_{V}𝟙, X), Λ^{𝔭}X ≅ Hom(Γ_{𝔭}𝟙, X) .

Proof. Combine the isomorphisms in (8.2) with the adjunction defining Hom. □

2The exactness of Hom(−, Y) was omitted from [4, §8], since it was not used there, but it is important in Lemma 8.4 below.

3For these results to hold, the R action should be canonical, for the R-linearity of the adjunction isomorphism (8.1) is used in the arguments.

Colocalizing subcategories. Function objects turn localizing subcategories into colocalizing subcategories in the following sense.

Lemma 8.4. Let C be a class of objects in T and X, Y ∈ T. If X belongs to Loc(C), then Hom(X, Y) belongs to Coloc({Hom(C, Y) | C ∈ C}).

Proof. This holds as Hom(−, Y) is exact and turns coproducts into products. □

We focus attention on colocalizing subcategories satisfying the equivalent conditions of the following lemma.

Lemma 8.5. Let S be a colocalizing subcategory of T. Then the following conditions on S are equivalent:

(1) For all compact objects X in T and all Y in S, X ⊗ Y is also in S.

(2) For all compact objects X in T and all Y in S, Hom(X, Y) is in S.

(3) For all objects X in T and all Y in S, Hom(X, Y) is in S.

Proof. The equivalence of (1) and (2) follows from the isomorphisms X^{∨} ⊗ Y ≅ Hom(X, Y) and X^{∨∨} ≅ X. The equivalence of (2) and (3) follows from Lemma 8.4 and the fact that T is compactly generated. □

We say that a colocalizing subcategory is Hom closed if the equivalent conditions of the lemma hold, and write Coloc^{Hom}(C) for the smallest Hom closed colocalizing subcategory containing a class C of objects in T.

A localizing subcategory S of T is tensor ideal or tensor closed if X ∈ T and Y ∈ S imply that X ⊗ Y is in S. Given a class C of objects in T, we denote by Loc^{⊗}(C) the smallest tensor ideal localizing subcategory containing C.

The gist of the next result is that (an appropriate version of) the local-global principle holds for tensor triangulated categories. The first part, about localizing subcategories, is from [5, Theorem 3.6].

Theorem 8.6. For each X ∈ T, there are equalities

Loc_{T}^{⊗}(X) = Loc_{T}^{⊗}({Γ_{𝔭}X | 𝔭 ∈ Spec R})
Coloc_{T}^{Hom}(X) = Coloc_{T}^{Hom}({Λ^{𝔭}X | 𝔭 ∈ Spec R}) .

Proof. The first equality is [5, Theorem 3.6].

In particular, Loc_{T}^{⊗}(𝟙) = Loc_{T}^{⊗}({Γ_{𝔭}𝟙 | 𝔭 ∈ Spec R}), which in conjunction with Lemma 8.4 gives the second equality below:

Coloc_{T}^{Hom}(X) = Coloc_{T}^{Hom}(Hom(𝟙, X))
= Coloc_{T}^{Hom}({Hom(Γ_{𝔭}𝟙, X) | 𝔭 ∈ Spec R})
= Coloc_{T}^{Hom}({Λ^{𝔭}X | 𝔭 ∈ Spec R})

The first one holds because X = Hom(𝟙, X) and last one is by Proposition 8.3. □

Lemma 8.7. If S is a colocalizing subcategory of T then Hom(X, Y) is in S for all X in Loc_{T}(𝟙) and Y in S. In particular, if 𝟙 generates T then all its colocalizing subcategories are Hom closed.

Proof. This follows from Lemma 8.4. □

Remark 8.8. Let T be a tensor triangulated category with a canonical R-action. If T is generated by its unit 𝟙, then the local global principle for (co)localizing subcategories holds. This follows from Theorem 8.6 and Lemma 8.7.

9. Costratification

Let T = (T, ⊗, 𝟙) be a tensor triangulated category as in Section 8, endowed with a canonical R-action. In this section, we introduce a variant of the notion of costratification, see Remark 5.7, suitable for this context and explain some consequences, including a classification of the Hom closed colocalizing subcategories.

Recall from Proposition 5.6 that there are maps σ and τ which yield a classification of colocalizing subcategories. Proposition 8.3 implies that each Λ^{𝔭}T is Hom closed, so these maps restrict to the following maps on Hom closed subcategories:

\{ \begin{matrix}Hom closed colocalizing \\ subcategories of T \end{matrix} \} [figure: σ; τ]\{ \begin{matrix}families (S(𝔭))_{𝔭∈Spec R} with \\ S(𝔭) ⊆ Λ^{𝔭}T a Hom closed \\ colocalizing subcategory \end{matrix} \}

where σ(S) = (S∩Λ^{𝔭}T) and τ(S(𝔭)) is the colocalizing subcategory of T cogenerated by all the S(𝔭). The following result is the analogue of [6, Proposition 3.6].

Proposition 9.1. The maps σ and τ are mutually inverse bijections.

Proof. The proof is exactly analogous to that of Proposition 5.6, using the localglobal principle from Theorem 8.6. □

We say that the tensor triangulated category T is costratified by R if for each 𝔭 ∈ Spec R, the colocalizing subcategory Λ^{𝔭}T contains no proper non-zero Hom closed colocalizing subcategories. Compare this definition with the one in Remark 5.7 for general triangulated categories. One does not have to impose the analogue of the local-global principle (C1), for it always holds; see Theorem 8.6.

If T is costratified by R then for each 𝔭 ∈ Spec R there are only two possibilities for S(𝔭), namely S(𝔭) = Λ^{𝔭}T or S(𝔭) = 0. So the maps σ and τ reduce to

\{ \begin{matrix}Hom closed colocalizing \\ subcategories of T \end{matrix} \} [figure: ^{cosupp^{R}}; _{cosupp{−1 \atop R}}]\{ subsets of supp_{R} T \} .

The next result is now immediate from the definition of costratification.

Corollary 9.2. If the tensor triangulated category T is costratified by R then the above maps cosupp_{R} and cosupp_{R}^{−1} are mutually inverse bijections. □

Stratification. In analogy with the notion of costratification, the tensor triangulated category T is stratified by R if for each 𝔭 ∈ Spec R the localizing subcategory Γ_{𝔭}T is zero or minimal among tensor ideal localizing subcategories; see [6, §7].

Next we establish a formula relating support and cosupport when T is stratified.

Lemma 9.3. An inclusion cosupp_{R} Hom(X, Y) ⊆ supp_{R} X ∩ cosupp_{R} Y holds for all objects X and Y in T.

Proof. Fix 𝔭 ∈ Spec R; recall that 𝔭 is in cosupp_{R} X if and only if Hom(Γ_{𝔭}𝟙, X) ≠ 0. Using (8.1) and (8.2), and Proposition 8.3, one gets isomorphisms

(9.4)  Hom(Γ_{𝔭}𝟙, Hom(X, Y)) ≅ Hom(Γ_{𝔭}X, Y) ≅ Hom(X, Λ^{𝔭}Y) .

It follows that Γ_{𝔭}X ≠ 0 and Λ^{𝔭}Y ≠ 0 when Λ^{𝔭} Hom(X, Y) ≠ 0. □

In [6, Theorem 7.3] we proved that supp_{R}(X ⊗ Y) = supp_{R} X ∩ supp_{R} Y holds if T is stratified. An analogue for function objects is contained in the next result.

Theorem 9.5. The following conditions on T are equivalent:

(1) The tensor triangulated category T is stratified by R.

(2) cosupp_{R} Hom(X, Y) = supp_{R} X ∩ cosupp_{R} Y for all X, Y ∈ T.
(3) Hom(X, Y) = 0 implies supp_{R} X ∩ cosupp_{R} Y = ∅ for all X, Y ∈ T.

Proof. (1) ⇒ (2): One inclusion follows from Lemma 9.3. For the other inclusion one uses that T is stratified by R. Fix 𝔭 ∈ supp_{R} X ∩ cosupp_{R} Y. The minimality of the tensor ideal localizing subcategory Γ_{𝔭}T implies Γ_{𝔭}𝟙 ∈ Loc^{⊗}(Γ_{𝔭}X), since Γ_{𝔭}X ≠ 0. Applying Lemma 8.4, one obtains

0 ≠ Hom(Γ_{𝔭}𝟙, Y) ∈ Coloc^{Hom}(Hom(Γ_{𝔭}X, Y))

and therefore Hom(Γ_{𝔭}X, Y) ≠ 0. Using the first isomorphism in (9.4), it follows that 𝔭 is in the cosupport of Hom(X, Y).

(2) ⇒ (3): Clearly, Hom(X, Y) = 0 implies cosupp_{R} Hom(X, Y) = ∅.

(3) ⇒ (1): To prove that the tensor triangulated category T is stratified by R, it suffices to show that given non-zero objects X and Y in Γ_{𝔭}T for some 𝔭 in Spec R, there exists a Z such that Hom_{T}^{∗}(X ⊗ Z, Y) ≠ 0; see [4, Lemma 3.9].

Since supp_{R} Y = {𝔭}, it follows from Theorem 4.13 that 𝔭 ∈ cosupp_{R} Y holds, and hence from our assumption that Hom(X, Y) ≠ 0. In particular, there exists a Z ∈ T such that Hom_{T}^{∗}(Z, Hom(X, Y)) ≠ 0. The adjunction isomorphism (8.1) then yields Hom_{T}^{∗}(X ⊗ Z, Y) ≠ 0. □

The preceding result has the following immediate consequence.

Corollary 9.6. Suppose the tensor triangulated category T is generated by its unit. Then T is stratified by R if and only if for all objects X and Y in T one has

Hom_{T}^{∗}(X, Y) = 0 ⟺ supp_{R} X ∩ cosupp_{R} Y = ∅.  □

There is the following connection between stratification and costratification.

Theorem 9.7. When the tensor triangulated category T is costratified by R, it is also stratified by R, and then there is an equality

cosupp_{R} Hom(X, Y) = supp_{R} X ∩ cosupp_{R} Y for all X, Y ∈ T.

Proof. It suffices to prove that given non-zero objects X and Y in Γ_{𝔭}T for some 𝔭 in Spec R, there exists a Z such that Hom_{T}^{∗}(X ⊗ Z, Y) ≠ 0; see [4, Lemma 3.9].

Assume T is costratified by R. As Γ_{𝔭}X ≠ 0 there exist an object C ∈ T such that Hom_{T}^{∗}(X, C(𝔭)) ≠ 0, by Proposition 4.14. It is easy to verify using the adjunction isomorphism (8.1) that the subcategory

S = {W ∈ T | Hom_{T}^{∗}(X ⊗ Z, W) = 0 for all Z ∈ T}

of T is colocalizing and Hom closed.

Now observe that 𝔭 ∈ cosupp_{R} Y holds by Theorem 4.13, since supp_{R} Y = {𝔭}. Thus Λ^{𝔭}T = Coloc^{Hom}(Λ^{𝔭}Y), by the costratification hypothesis. This implies Λ^{𝔭}Y ∉ S, since C(𝔭) ∈ Λ^{𝔭}T by Theorem 4.13, and C(𝔭) ∉ S. Thus one obtains that

Hom_{T}^{∗}(X ⊗ Z, Y) ≅ Hom_{T}^{∗}(Γ_{𝔭}(X ⊗ Z), Y) ≅ Hom_{T}^{∗}(X ⊗ Z, Λ^{𝔭}Y) ≠ 0

for some Z in T, and hence that T is stratified.

The formula for cosupp_{R} Hom(X, Y) follows from Theorem 9.5. □

Remark 9.8. It is an open question whether stratification implies costratification. The proof of Theorem 9.7 uses the fact that every localizing subcategory generated by a set of objects arises as the kernel of a localization functor; see the proof of [4, Lemma 3.9]. It is not known whether the analogous statement for colocalizing subcategories is true or not. This reflects the fact that products are usually more complicated than coproducts.

The following corollary combines the classification of colocalizing subcategories, Corollary 9.2, with the classification of localizing subcategories in [5, Theorem 3.8].

Corollary 9.9. If the tensor triangulated category T is costratified by R then the map sending a subcategory S to S^{⊥} induces a bijection

\{ \begin{matrix}tensor closed localizing \\ subcategories of T \end{matrix} \} [figure: ∼]\{ \begin{matrix}Hom closed colocalizing \\ subcategories of T \end{matrix} \} .

The inverse map sends a Hom closed colocalizing subcategory U to ^{⊥}U.

Proof. Assume T is costratified by R; it is then stratified by R, by Theorem 9.7. In particular, both the tensor closed localizing subcategories and the Hom closed colocalizing subcategories of T are in bijection with the subsets of supp_{R} T, via the maps supp_{R}(−) and cosupp_{R}(−), respectively; see [4, Theorem 3.8] and Corollary 9.2. Now, for any tensor closed localizing subcategory S of T one has equalities

S^{⊥} = {Y ∈ T | Hom(X, Y) = 0 for all X ∈ S}
= {Y ∈ T | cosupp_{R} Y ∩ supp_{R} S = ∅}
= cosupp_{R}^{−1}(supp_{R} T ∖ supp_{R} S)

where the first one is a routine verification, and the second one is by Theorem 9.7. In the same vein, for any Hom closed localizing subcategory U one has

⊥U = supp_{R}^{−1}(supp_{R} T ∖ cosupp_{R} U).

It thus follows that, under the identification above, both maps S ↦ S^{⊥} and U ↦ ^{⊥}U correspond to the map on supp_{R} T sending a subset to its complement, and are thus inverse to each other. □

Brown–Comenetz duality. Let k be a commutative ring and suppose that the category T is k-linear. We denote by D = Hom_{k}(−, E) the duality for the category of k-modules with respect to a fixed injective cogenerator E.

The Brown–Comenetz dual X^{∗} of an object X is defined by the isomorphism

D Hom_{T}(𝟙, − ⊗ X) ≅ Hom_{T}(−, X^{∗}).

Note that there is a natural isomorphism X^{∗} ≅ Hom(X, 𝟙^{∗}).

Proposition 9.10. One has cosupp_{R} 𝟙^{∗} = supp_{R} T and cosupp_{R} X^{∗} ⊆ supp_{R} X for any X ∈ T; equality holds if T is stratified by R as a tensor triangulated category.

Proof. The first equality holds because X^{∗} ≅ Hom(X, 𝟙^{∗}) and X^{∗} = 0 if and only if X = 0. The inclusion follows from Lemma 9.3, since

cosupp_{R} X^{∗} = cosupp_{R} Hom(X, 𝟙^{∗}) ⊆ supp_{R} X ∩ cosupp_{R} 𝟙^{∗} = supp_{R} X.

When T is stratified, Theorem 9.5 gives equality. □

Remark 9.11. Let T be a tensor triangulated category generated by its unit 𝟙, and equipped with a canonical R action. Given Remark 8.8, it is clear that T is (co)stratified by R as a triangulated category if, and only if, it is (co)stratified by R as a tensor triangulated category. And to verify that T is stratified (respectively, costratified) it suffices to check that each Γ_{𝔭}T is a minimal localizing subcategory (respectively, each Λ^{𝔭}T is a minimal colocalizing subcategory).

The results in this section thus yield interesting information about all localizing and colocalizing subcategories of T. For instance, Corollary 9.2 coincides with the bijection in Remark 5.7, while Theorem 9.7 says that if T is costratified then it is also stratified.

10. Formal dg algebras

In this section we prove that the derived category of dg (short for “differential graded”) modules over a formal dg algebra is costratified by its cohomology algebra, when that algebra is graded commutative and noetherian.

Let A be a dg algebra and D(A) its derived category of (left) dg modules. It is a triangulated category, generated by the compact object A; see [16]. A morphism A → B of dg algebras is a quasi-isomorphism if the induced map H^{∗}A → H^{∗}B is an isomorphism. Then restriction induces an equivalence of triangulated categories D(B) ⟶^{∼} D(A), with quasi-inverse the functor B ⊗_{A}^{L} −. Dg algebras A and B are called quasi-isomorphic if there is a finite chain of quasi-isomorphisms linking them.

The multiplication on A induces one on its cohomology H^{∗}A. We say A is formal if it is quasi-isomorphic to H^{∗}A, viewed as a dg algebra with zero differential.

Remark 10.1. Suppose that A is formal and that H^{∗}A is graded commutative. Fix a chain of quasi-isomorphisms linking A and H^{∗}A; it induces an equivalence of categories D(A) ≃ D(H^{∗}A).

The derived tensor product of dg modules endows D(H^{∗}A) with a structure of a tensor triangulated category, with unit H^{∗}A. Thus, D(A) also acquires such a structure, via the equivalence D(A) ≃ D(H^{∗}A); denote ⊗ this tensor product on D(A). The object A is a tensor unit, and one gets an action of H^{∗}A on D(A), defined by taking for each object X the composite map

H^{∗}A ⟶^{∼} End_{D(A)}^{∗}(A) ⟶^{X⊗−} End_{D(A)}^{∗}(X) .

We refer to this as the action induced by the given chain of quasi-isomorphisms linking A and H^{∗}A.

The action of H^{∗}A on D(A) depends on the choice of quasi-isomorphisms linking A and H^{∗}A. One has however the following independence statement.

Lemma 10.2. Let A be a formal dg algebra with H^{∗}A graded commutative and noetherian. For any specialization closed set V ⊆ Spec H^{∗}A the functors Γ_{V}, L_{V}, Λ^{V} and V^{V} are independent of a chain of quasi-isomorphisms linking A to H^{∗}A.

Proof. For any chain of quasi-isomorphisms, it is clear from the construction that the homomorphism H^{∗}A → End_{D(A)}^{∗}(A) is the canonical one, and hence it is independent of the action. It thus remains to note that the local cohomology functors on a compactly generated R-linear triangulated category are determined by the action of R on a compact generator, by [5, Corollary 3.3]. □

In the case when A is a ring, which may be viewed as a dg algebra concentrated in degree zero, the result below is contained in recent work of Neeman [21]. We note that the cosupport, and hence the costratification, are independent of a choice of a canonical action, by Lemma 10.2.

Theorem 10.3. Let A be a formal dg algebra such that H^{∗}A is graded commutative and noetherian. The category D(A) is costratified by any H^{∗}A-action induced by a chain of quasi-isomorphisms linking A and H^{∗}A.

Proof. We may replace A by H^{∗}A and assume d^{A} = 0 and A is graded commutative. Set D = D(A). Since A is a unit and a generator of this tensor triangulated category, its colocalizing subcategories are RHom closed; see Lemma 8.7. It remains to verify that Λ^{𝔭}D is a minimal colocalizing subcategory for each homogeneous prime ideal 𝔭 in A; see Remark 9.11.

Let k(𝔭) be the graded residue field at 𝔭. The object k(𝔭) is in Λ^{𝔭}D, so to verify the minimality of Λ^{𝔭}D it suffices to note that for each non-zero M in Λ^{𝔭}D, the following equalities hold:

Coloc_{D}(M) = Coloc_{D}(RHom_{A}(Γ_{𝔭}A, M))
= Coloc_{D}(RHom_{A}(k(𝔭), M))
= Coloc_{D}(k(𝔭))

The first equality holds by Proposition 8.3 since M ≅ Λ^{𝔭}M. As to the second one, since A is stratified by H^{∗}A, by [6, Theorem 8.1], one has

Loc_{D}(Γ_{𝔭}A) = Γ_{𝔭}D = Loc_{D}(k(𝔭)) .

Now apply Lemma 8.4. The last equality follows from the fact that k(𝔭) is a graded field and the action of A on RHom_{A}(k(𝔭), M) factors through k(𝔭). □

Exterior algebras. Let Λ be a graded exterior algebra over a field k on indeterminates ξ_{1}, …, ξ_{c} in negative odd degree, regarded as a dg algebra with zero differential. We give Λ a structure of Hopf algebra via Δ(ξ_{i}) = ξ_{i} ⊗ 1 + 1 ⊗ ξ_{i}. In [5, §4], we introduced the homotopy category of graded-injective dg modules over Λ, denoted K(Inj Λ), and proved the following statements: K(Inj Λ) is a compactly generated tensor triangulated category, in the sense of Section 8; its unit is an injective resolution of the trivial module k, and this generates K(Inj Λ); the graded endomorphism algebra of the unit is Ext_{Λ}^{∗}(k, k), which is isomorphic to the graded polynomial k-algebra S = k[x_{1}, …, x_{c}] with |x_{i}| = −|ξ_{i}| + 1.

Theorem 10.4. The category K(Inj Λ) is costratified by the action of Ext_{Λ}^{∗}(k, k).

Proof. It is proved in [5, Theorem 6.2] that a suitable dg module J over Λ ⊗_{k} S yields an equivalence Hom_{Λ}(J, −): K(Inj Λ) ⟶^{∼} D(S). The theorem now follows from Theorem 10.3. □

11. Finite groups

Throughout this section, G will be a finite group and k a field whose characteristic divides the order of G. The associated group algebra is denoted kG, and H^{∗}(G, k) denotes its cohomology k-algebra, Ext_{kG}^{∗}(k, k). This algebra is graded commutative, because kG is a Hopf algebra, and finitely generated and hence noetherian, by a result of Evens and Venkov [10, 24]; see also Golod [13].

Let K(Inj kG) be the homotopy category of complexes of injective kG-modules. This is a compactly generated tensor triangulated category, in the sense of Section 8, where the tensor product X ⊗_{k} Y and the function object Hom_{k}(X, Y) are induced by those on kG-modules via the diagonal action of G. The injective resolution ik of the trivial module k is the identity for the tensor product, and it yields a canonical action of H^{∗}(G, k) on K(Inj kG); see [7] for details. To simplify notation, we write V_{G} for Spec H^{∗}(G, k), and set for each object X in K(Inj kG)

supp_{G} X = supp_{H^{∗}(G,k)} X and cosupp_{G} X = cosupp_{H^{∗}(G,k)} X.

Example 11.1. We write 𝔪 for the maximal ideal H^{⩾1}(G, k).

(1) One has cosupp_{G} X = {𝔪} for any non-zero compact object X ∈ K(Inj kG).

(2) For each 𝔭 ∈ V_{G} the object T_{ik}(I(𝔭)), from (5.3), satisfies

supp_{G} T_{ik}(I(𝔭)) = {𝔭} and cosupp_{G} T_{ik}(I(𝔭)) = {𝔮 ∈ V_{G} | 𝔮 ⊆ 𝔭} .

Indeed, for (1) observe that the natural morphism X → X^{∗∗} is an isomorphism for each compact object X by [5, Lemma 11.5], where (−)^{∗} denotes Brown– Comenetz duality. Proposition 9.10 therefore gives the first inclusion below

cosupp_{G} X ⊆ supp_{G} X^{∗} ⊆ {𝔪} .

For the second inclusion one uses that the H^{∗}(G, k)-module Hom_{K(kG)}^{∗}(C, X^{∗}) is 𝔪-torsion for each compact C ∈ K(Inj kG), since it is of the form Hom_{k}(M, k) for some finitely generated H^{∗}(G, k)-module M by the defining isomorphism of the Brown–Comenetz dual X^{∗}. It remains to observe that cosupp_{G} X ≠ ∅ since X ≠ 0.

To prove (2), note that Hom_{K(kG)}^{∗}(C, T_{ik}(I(𝔭))) is 𝔭-torsion and 𝔭-local for each compact object C ∈ K(Inj kG), by the isomorphism (5.3) defining T_{ik}(I(𝔭)). Thus supp_{G} T_{ik}(I(𝔭)) = {𝔭}, while the cosupport is given by Proposition 5.4.

Restriction and induction. For each subgroup H of G restriction and induction yield exact functors

(−)↓_{H} = Hom_{kG}(kG, −): K(Inj kG) → K(Inj kH) and
(−)↑^{G} = − ⊗_{kH} kG: K(Inj kH) → K(Inj kH) .

Restriction yields also a homomorphism res_{G,H}: H^{∗}(G, k) → H^{∗}(H, k) of graded rings, and hence a map:

res_{G,H}^{∗}: V_{H} → V_{G} .

It is easy to verify that these functors fit in the framework of Section 7:

Lemma 11.2. The functor (−)↓_{H} is res_{G,H}-linear and induces an exact functor

((−)↓_{H}, res_{G,H}): (K(Inj kG), H^{∗}(G, k)) → (K(Inj kH), H^{∗}(H, k))

which preserves coproducts and products. Its right adjoint and left adjoint is (−)↑^{G}, and the latter is faithful on objects. □

The following result is an analogue of [5, Lemma 9.3].

Lemma 11.3. Let H be a subgroup of G. Fix 𝔭 ∈ V_{G} and set U = (res_{G,H}^{∗})^{−1}{𝔭}. The set U is finite and discrete, and for any X ∈ K(Inj kG) and any Y ∈ K(Inj kH) there are natural isomorphisms:

(Γ_{𝔭}X)↓_{H} ≅ ∐_{𝔮∈U} Γ_{𝔮}(X↓_{H}) Γ_{𝔭}(Y ↑^{G}) ≅ ∐_{𝔮∈U}(Γ_{𝔮}Y)↑^{G}

(Λ^{𝔭}X)↓_{H} ≅ ∏_{𝔮∈U} Λ^{𝔮}(X↓_{H}) Λ^{𝔭}(Y ↑^{G}) ≅ ∏_{𝔮∈U}(Λ^{𝔮}Y)↑^{G}

Proof. The set U is finite and discrete as H^{∗}(H, k) is finitely generated as a module over H^{∗}(G, k). The result now follows from Corollary 7.10, given Lemma 11.2. □

The next result is an analogue of [5, Proposition 9.4].

Proposition 11.4. Let H be a subgroup of G. The following statements hold: (1) For any object Y in K(Inj kH), there is an equality

cosupp_{G}(Y ↑^{G}) = res_{G,H}^{∗}(cosupp_{H} Y) .

(2) Any object X in K(Inj kG) satisfies X↓_{H}↑^{G} ∈ Coloc^{Hom}(X), and hence

cosupp_{G}(X↓_{H}↑^{G}) ⊆ cosupp_{G} X .

Proof. Part (1) is a special case of Corollary 7.8, which applies by Lemma 11.2.

(2) One has X↓_{H}↑^{G} ≅ X ⊗_{k} k(G/H), where k(G/H) is the permutation module on the cosets of H in G. Let W be an injective resolution over kG of k(G/H). Arguing as in the proof of [7, Proposition 5.3], one can prove that the natural map X ⊗_{k} k(G/H) → X ⊗_{k} W is an isomorphism in K(Inj kG). As W is compact in K(Inj kG), see [18, Lemma 2.1], it remains to apply Lemma 8.5. □

Elementary abelian groups. Let E be an elementary abelian p-group

E = ⟨g_{1}, …, g_{r}⟩

and k a field of characteristic p. We set z_{i} = g_{i} − 1, so that the group algebra may be described as

kE = k[z_{1}, …, z_{r}]/(z_{1}^{p}, …, z_{r}^{p}) .

Let A be the Koszul complex on z_{1}, …, z_{r}, viewed as a dg algebra: The graded algebra underlying it is generated by kE in degree zero together with exterior generators y_{1}, …, y_{r}, each of degree −1. The differential on A is given by d(y_{i}) = z_{i} and d(z_{i}) = 0. We write K(Inj A) for the homotopy category of graded-injective dg A-modules. It is a compactly generated tensor triangulated category, with a canonical action of Ext_{A}^{∗}(k, k); see [5, §8].

Proposition 11.5. The category K(Inj A) is costratified by the action of Ext_{A}^{∗}(k, k).

Proof. Let Λ be an exterior algebra over k on indeterminates ξ_{1}, …, ξ_{r} of degree −1, viewed as a dg algebra with d^{Λ} = 0. Let K(Inj Λ) be the homotopy category of graded-injective dg Λ-modules, with tensor triangulated structure described in the paragraph preceding Theorem 10.4.

By [5, Lemma 7.1] there is a quasi-isomorphism of dg k-algebras ϕ: Λ → A defined by ϕ(ξ_{i}) = z_{i}^{p−1}y_{i}, and by [5, Proposition 4.6] this induces an equivalence of triangulated categories

Hom_{Λ}(A, −): K(Inj Λ) ⟶^{∼} K(Inj A) .

The desired result is now a consequence of Theorem 10.4. □

The next result complements [5, Theorem 8.1], concerning stratification.

Theorem 11.6. Let E be an elementary abelian p-group and k a field of characteristic p. The category K(Inj kE) is costratified by the canonical action of H^{∗}(E, k).

Proof. Write kE ≅ k[z_{1}, …, z_{r}]/(z_{1}^{p}, …, z_{r}^{p}), and let A be the Koszul dg algebra described above. Note that kE = A_{0} so the inclusion kE → A is a morphism of dg algebras; restriction along it gives an exact functor K(Inj A) → K(Inj kE) which preserves coproducts and products, and is compatible with the induced homomorphism α: Ext_{A}^{∗}(k, k) → Ext_{kE}^{∗}(k, k) = H^{∗}(E, k). The functor Hom_{kE}(A, −) is a right adjoint of restriction. We claim that each X ∈ K(Inj kE) satisfies an equality:

(∗)  Thick(Hom_{kE}(A, X)) = Thick(X) in K(Inj kE),

and, in particular, that Hom_{kE}(A, −) is faithful on objects.

Indeed, recall that A is the Koszul complex over kE on z = z_{1}, …, z_{r}, and in particular, a finite free complex of kE-modules. This yields the first isomorphism below of complexes of kE-modules:

Hom_{kE}(A, X) ≅ Hom_{kE}(A, kE) ⊗_{kE} X ≅ Σ^{r}A ⊗_{kE} X

The second one is by self-duality of the Koszul complex; see [8, Proposition 1.6.10]. The radical of the ideal (z) coincides with that of (0), so the complexes A and kE generate the same thick subcategory in K(Inj kE); see [15, Lemma 6.0.9]. Since − ⊗_{kE} X is an exact functor on K(Inj kE), it follows from the isomorphism above that Hom_{kE}(A, X) and X generate the same thick subcategory, as desired.

We are thus in a position to apply Corollary 7.8. Fix a prime 𝔭 ∈ V_{E} and let 𝔮 = α^{−1}(𝔭). The functor Hom_{kE}(A, −) then takes Λ^{𝔭}K(Inj kE) to Λ^{𝔮}K(Inj A) and is faithful. Hence, for any non-zero objects X and Y in Λ^{𝔭}K(Inj kE), one has

Hom_{kE}(A, Y) ∈ Coloc(Hom_{kE}(A, X)) in K(Inj A),

since K(Inj A) is costratified by the action of Ext_{A}^{∗}(k, k); see Proposition 11.5. It implies, in view of the equality (∗) above, that in K(Inj kE) one has

Y ∈ Thick(Hom_{kE}(A, Y)) ⊆ Coloc(Hom_{kE}(A, X)) = Coloc(X).

Thus, Λ^{𝔭}K(Inj kE) is a minimal colocalizing subcategory. Thus, K(Inj kE) is costratified by the action of H^{∗}(E, k); see Remark 9.11. □

Next we prepare for a version of the preceding theorem for arbitrary finite groups.

Quillen’s stratification. We consider pairs (H, 𝔮) of subgroups H of G and primes 𝔮 ∈ V_{H}, and say that (H, 𝔮) and (H′, 𝔮′) are G-conjugate if conjugation with some element in G takes H to H′ and 𝔮 to 𝔮′.

Lemma 11.7. Let (H, 𝔮) and (H′, 𝔮′) be G-conjugate pairs and X in K(Inj kG). Then Λ^{𝔮}(X↓_{H}) ≠ 0 if and only if Λ^{𝔮′}(X↓_{H′}) ≠ 0.

Proof. Conjugation with any element g ∈ G induces an automorphism of K(Inj kG) that takes an object X to X^{g}. Note that multiplication with g induces an isomorphism X ⟶^{∼} X^{g}. The assertion follows, since conjugation commutes with restriction and local homology. □

Quillen has proved that for each 𝔭 in V_{G} there exists a pair (E, 𝔮) such that E is an elementary abelian subgroup of G and res_{G,E}^{∗}(𝔮) = 𝔭; see the discussion after [22, Proposition 11.2]. We say 𝔭 originates in such a pair (E, 𝔮) if there does not exist another such pair (E′, 𝔮′) with E′ a proper subgroup of E. In this language, [22, Theorem 10.2] reads:

Theorem 11.8. For any 𝔭 ∈ V_{G} all pairs (E, 𝔮) where 𝔭 originates are G-conjugate. This sets up a one to one correspondence between primes 𝔭 in V_{G} and G-conjugacy classes of such pairs (E, 𝔮). □

The proof of the next result can be shortened considerably by invoking the subgroup theorem for supports, [4, Theorem 11.2], which is deduced from the stratification theorem for K(Inj kG), [4, Theorem 9.7]. We give a direct proof, by extracting an argument from the proof of the latter result.

Proposition 11.9. Fix 𝔭 ∈ V_{G} and suppose 𝔭 originates in E. Given a non-zero object X in Λ^{𝔭}K(Inj kG), one has

X↓_{E} = ∏_{res_{G,E}^{∗}(𝔮)=𝔭} Λ^{𝔮}(X↓_{E})

and Λ^{𝔮}(X↓_{E}) ≠ 0 for each such 𝔮.

Proof. The decomposition of X↓_{E} is by Lemma 11.3. For the remaining statements, it suffices to find one pair (E′, 𝔮′) where 𝔭 originates and such that Λ^{𝔮′}(X↓_{E′}) ≠ 0. Then Theorem 11.8 yields that each pair (E, 𝔮) as in the statement is G-conjugate to (E′, 𝔮′), and hence Λ^{𝔮}(X↓_{E}) ≠ 0, by Lemma 11.7.

Since X ≠ 0 holds, Chouinard’s theorem for K(Inj kG), see [4, Proposition 9.6(3)], provides an elementary abelian subgroup E′′ of G such that X↓_{E′′} ≠ 0. Fix a 𝔮′′ in cosupp_{E′′}(X↓_{E′′}). Proposition 11.4 then yields

res_{G,E′′}^{∗}(𝔮′′) ∈ cosupp_{G}(X↓_{E′′}↑^{G}) ⊆ cosupp_{G} X = {𝔭} .

Applying Theorem 11.8 to E′′ one gets a pair (E′, 𝔮′), with E′ a subgroup of E′′, where 𝔮′′ originates. Observe that 𝔭 then originates in (E′, 𝔮′), by functoriality of restrictions and the computation above. It remains to note that

𝔮′ ∈ (res_{E′′,E′}^{∗})^{−1}(cosupp_{E′′}(X↓_{E′′})) = cosupp_{E′}(X↓_{E′}) ,

where the inclusion holds by the choice of 𝔮′ and the equality is by Theorem 7.11, applied to the functor

((−)↓_{E′}, res_{E′′,E′}): (K(Inj kE′′), H^{∗}(E′′, k)) → (K(Inj kE′), H^{∗}(E′, k)) ;

noting that the hypotheses are satisfied by Lemma 11.2 and Theorem 11.6. □

Finite groups. The following theorem is an analogue of [5, Theorem 9.7], which establishes the stratification of K(Inj kG); it contains Theorem 11.6, but the latter statement is used in the proof, both directly and by way of Proposition 11.9.

Theorem 11.10. The tensor triangulated category K(Inj kG) is costratified by the canonical action of the cohomology algebra H^{∗}(G, k).

Proof. Fix 𝔭 ∈ V_{G} and suppose it originates in E. For each compact object C in K(Inj kE) and each 𝔮 ∈ V_{E} with res_{G,E}^{∗}(𝔮) = 𝔭, Proposition 5.4 and (3.1) imply

cosupp_{E}(T_{C//𝔮}(I(𝔮))) ⊆ {𝔮} .

It follows from Theorem 11.6 and Proposition 11.9 that each non-zero object X in Λ^{𝔭}K(Inj kG) satisfies

T_{C//𝔮}(I(𝔮)) ∈ Coloc(Λ^{𝔮}(X↓_{E})) ⊆ Coloc(X↓_{E}) .

Together with Proposition 11.4, one thus obtains:

T_{C//𝔮}(I(𝔮))↑^{G} ∈ Coloc(X↓_{E}↑^{G}) ⊆ Coloc^{Hom}(X) .

It remains to observe that the collection of objects T_{C//𝔮}(I(𝔮))↑^{G} cogenerates the triangulated category Λ^{𝔭}K(Inj kG). This is a consequence of Proposition 5.4 and Lemma 7.12, which can be applied, thanks to Lemma 11.2 and Proposition 11.9. □

Applications. The consequences of costratification described in Section 9 apply to K(Inj kG). In particular, Theorem 9.7 implies that K(Inj kG) is stratified as a tensor triangulated category by H^{∗}(G, k), which is [5, Theorem 9.7]. A modification of Theorem 7.11 yields the following subgroup theorem for cosupport, which is analogous to the one for support; see [5, Theorem 11.2].

Theorem 11.11. Let H be a subgroup of G. Each object X in K(Inj kG) satisfies

cosupp_{H}(X↓_{H}) = (res_{G,H}^{∗})^{−1}(cosupp_{G} X).

Proof. Restriction and induction form an adjoint pair of functors satisfying the assumptions of Theorem 7.11. Thus one gets

cosupp_{H}(X↓_{H}) ⊆ (res_{G,H}^{∗})^{−1}(cosupp_{G} X).

For the other inclusion, one uses stratification of K(Inj kG) as follows.

Fix a 𝔮 ∈ V_{H} with 𝔮 ∉ cosupp_{H}(X↓_{H}). We need to show that 𝔭 = res_{G,H}^{∗}(𝔮) is not in cosupp_{G} X. Using the adjunction formula for function objects from Proposition 8.3, one has

Hom_{k}(Γ_{𝔮}−, X↓_{H}) ≅ Hom_{k}(−, Λ^{𝔮}X↓_{H}) = 0.

This implies Hom_{k}((Γ_{𝔮}−)↑^{G}, X) = 0, since

Hom_{k}(U, V ↓_{H})↑^{G} ≅ Hom_{k}(U↑^{G}, V)

for all U in K(Inj kH) and V in K(Inj kG), by [1, Proposition 3.3]. There exists some U in K(Inj H) such that (Γ_{𝔮}U)↑^{G} ≠ 0 since induction is faithful on objects. Moreover, (Γ_{𝔮}U)↑^{G} belongs to Γ_{𝔭}K(Inj kG), by Corollary 7.8. Since K(Inj kG) is stratified as a tensor triangulated category by H^{∗}(G, k), the subcategory Γ_{𝔭}K(Inj kG) contains no non-trivial tensor ideal localizing subcategories, and hence coincides with Loc^{⊗}((Γ_{q}U)↑^{G}). Thus

0 = Hom_{k}(Γ_{𝔭}−, X) ≅ Hom_{k}(−, Λ^{𝔭}X),

again by Proposition 8.3, and therefore 𝔭 ∉ cosupp_{R} X. □

Stable module category. Let Mod(kG) denote the (abelian) category of all kGmodules. The stable module category, StMod(kG), has the same objects as the module category Mod(kG), but the morphisms in StMod(kG) are given by quotienting out those morphisms in Mod(kG) that factor through a projective module. The category StMod(kG) is tensor triangulated, with triangles induced from short exact sequences of kG-modules, and product the tensor product over k with diagonal action. The trivial module k is the unit of the product, and its graded endomorphism ring in StMod(kG) is the Tate cohomology algebra Ĥ^{∗}(G, k). There is thus an action of H^{∗}(G, k) on StMod(kG), via the natural map H^{∗}(G, k) → Ĥ^{∗}(G, k).

We regard StMod(kG) as a triangulated subcategory of K(Inj kG), as in [7, §6].

Proposition 11.12. The functor StMod(kG) → K(Inj kG) that takes a kG-module to its Tate resolution identifies the tensor triangulated category StMod(kG) with the (co)localizing subcategory consisting of all acyclic complexes. □

The action of H^{∗}(G, k) on StMod(kG) is compatible with this identification. Moreover the acyclic complexes form a localizing and colocalizing subcategory of K(Inj kG) that is Hom and tensor closed. Note that a complex is acyclic if and only if its (co)support does not contain the maximal ideal of H^{∗}(G, k). This follows from [5, Proposition 9.6] for the support, and then from Corollary 4.9 for the cosupport.

Theorem 11.13. The tensor triangulated category StMod(kG) is costratified by the canonical action of H^{∗}(G, k).

Proof. Given Proposition 11.12, the desired result follows from Theorem 11.10, since Λ^{𝔭} StMod(kG) = Λ^{𝔭}K(Inj kG) for each non-maximal 𝔭 in Spec H^{∗}(G, k). □

We can now justify the results stated in the introduction.

Proof of Theorem 1.1. From Theorems 11.13 and 9.7 it follows that the tensor triangulated category StMod(kG) is stratified and costratified by the canonical action of H^{∗}(G, k). Thus the map sending a subset U of V_{G} to the subcategory of kGmodules X satisfying supp_{G} X ⊆ U yields a bijection between subsets of V_{G} and tensor ideal localizing subcategories of StMod(kG); see [5, Theorem 3.8]. Composing this bijection with the one between localizing and colocalizing subcategories from Corollary 9.9 gives the desired result. □

Proof of Corollary 1.2. Given Theorem 11.13, this follows from Corollary 9.9. □

We close this discussion on the stable module category with the following example. There is no analogue for arbitrary tensor triangulated categories; see Example 11.1 and Proposition 4.19.

Example 11.14. There is an equality cosupp_{G} M = supp_{G} M for any finite dimensional module M in StMod(kG).

Indeed, denote by (−)^{∗} the Brown–Comenetz duality on StMod(kG) which equals Ω Hom_{k}(−, k) by [5, Proposition 11.6], where ΩN denotes the kernel of a projective cover of a kG-module N. Then one gets the first equality below because M is finite dimensional:

cosupp_{G} M = cosupp_{G} M^{∗∗} = supp_{G} M^{∗} = supp_{G} M.

The second equality is Proposition 9.10, and the last one is well-known; see [2, Theorem 5.1.1].

Modules. Although Mod(kG) is not a triangulated category, we define colocalizing subcategories in an analogous way. A full subcategory of Mod(kG) is said to be thick if whenever two modules in a short exact sequence are in, then so is the third. A colocalizing subcategory S of Mod(kG) is a thick subcategory closed under all products, that is, for any family of modules M_{i} (i ∈ I) in S the product ∏_{i} M_{i} is in S. The next lemma describes an extra tensor condition for colocalizing subcategories; it parallels Lemma 8.5.

Lemma 11.15. Let S be a colocalizing subcategory of Mod(kG). Then the following conditions are equivalent:

(1) If N is a simple kG-module and M is in S then N ⊗_{k} M is in S.

(2) If N is a finitely generated kG-module and M is in S then N ⊗_{k} M is in S.

(3) If N is a kG-module and M is in S then Hom_{k}(N, M) is in S. □

A colocalizing subcategory of Mod(kG) is said to be Hom closed if the equivalent conditions of the lemma hold. The next result is analogous to [5, Proposition 2.1].

Proposition 11.16. The canonical functor from Mod(kG) to StMod(kG) induces a one to one correspondence between non-zero Hom closed colocalizing subcategories of Mod(kG) and Hom closed colocalizing subcategories of StMod(kG). □

Thus the classification of Hom closed colocalizing subcategories of Mod(kG) is a consequence of the costratification of StMod(kG) formulated in Theorem 11.13.

Appendix A. Localization functors and their adjoints

Localization functors. A functor L: C → C is called localization functor if there exists a morphism η: Id_{C} → L such that the morphism Lη: L → L^{2} is invertible and Lη = ηL. The morphism η is called adjunction. Recall that a morphism η: F → F′ between functors is invertible if ηX: F X → F′X is invertible for each object X. A functor Γ: C → C is by definition a colocalization functor if its opposite functor Γ^{op} : C^{op} → C^{op} is a localization functor. The corresponding morphism Γ → Id_{C} is called coadjunction.

Any localization functor L: C → C can be written as the composite of two functors that form an adjoint pair. In order to define these functors let us denote by Im L the essential image of L, that is, the full subcategory of C that is formed by all objects isomorphic to one of the form LX for some X in C.

Lemma A.1. Let L: C → C be a localization functor and write L as a composite

C ⟶^{F} Im L ⟶^{G} C

where G denotes the inclusion functor. Then G is a right adjoint of F.

Proof. See [4, Lemma 3.1] □

Next we characterize the adjoint pairs of functors (F, G) such that its composite L = GF is a localization functor. Given any category C and a class Σ of morphisms in C, we denote by Q: C → C[Σ^{−1}] the universal functor that inverts all morphisms in Σ; see [12, §I.1]. Thus Qϕ is invertible for all ϕ in Σ, and any functor F: C → D such that F ϕ is invertible for all ϕ in Σ factors uniquely through Q.

Lemma A.2. Let F: C → D and G: D → C be a pair of functors such that G is a right adjoint of F. Let L = GF and η: Id_{C} → L be the adjunction morphism. Then the following are equivalent:

(1) The morphism Lη: L → L^{2} is invertible and Lη = ηL (that is, L is a localization functor).

(2) The functor F induces an equivalence C[Σ^{−1}] ⟶^{∼} D, where Σ denotes the class of morphisms ϕ in C such that F ϕ is invertible.

(3) The functor G is fully faithful.

Proof. For (1) ⇔ (3) see [4, Lemma 3.1]; for (2) ⇔ (3) see [12, Proposition I.1.3]. □

Let F: C → D be an exact functor between triangulated categories. Then the full subcategory Ker F consisting of all objects that are annihilated by F is a thick subcategory of C. Denote by Σ the class of morphisms ϕ in C such that F ϕ is invertible. Then a morphism in C belongs to Σ if and only if its cone belongs to Ker F. Thus C[Σ^{−1}] = C/ Ker F; see [25].

The right adjoint of a localization functor. We discuss the existence of a right adjoint of a localization functor.

Lemma A.3. Let L: C → C be a localization functor with a right adjoint Γ. Then Γ is a colocalization functor satisfying Im Γ = Im L. If η: Id_{C} → L denotes the adjunction of L, then the coadjunction θ: Γ → Id_{C} of Γ is given by the composite

Hom_{C}(X, Γ Y) ⟶^{∼} Hom_{C}(LX, Y) ⟶^{(ηX,Y)} Hom_{C}(X, Y), X, Y ∈ C .

Proof. Fix an object Y in C. We need to show that Γ (θY) is an isomorphism and that Γ (θY) = θ(Γ Y). Denote by

ϕ_{X,Y} : Hom_{C}(X, Γ Y) ⟶^{∼} Hom_{C}(LX, Y)

the natural bijection given by the adjunction between L and Γ. We use that L(ηX) is an isomorphism and that L(ηX) = η(LX). The following commutative diagram shows that Γ (θY) is an isomorphism since all horizontal maps are bijections.

\xymatrix{Hom_{C}(X, Γ^{2}Y) \ar[r]^{ϕ_{X,Γ Y}} \ar[d]_{(X,Γ θY)} & Hom_{C}(LX, Γ Y) \ar[r]^{ϕ_{LX,Y}} \ar[d]^{(LX,θY)} & Hom_{C}(L^{2}X, Y) \ar[d]^{(ηLX,Y)} \\ Hom_{C}(X, Γ Y) \ar[r]^{ϕ_{X,Y}} & Hom_{C}(LX, Y) \ar[r]^{id} & Hom_{C}(LX, Y)}

Combining the previous diagram with the next one shows that Γ (θY) = θ(Γ Y).

\xymatrix{Hom_{C}(X, Γ^{2}Y) \ar[r]^{ϕ_{X,Γ Y}} \ar[d]_{(X,θΓ Y)} & Hom_{C}(LX, Γ Y) \ar[r]^{ϕ_{LX,Y}} \ar[d]^{(ηX,Γ Y)} & Hom_{C}(L^{2}X, Y) \ar[d]^{(LηX,Y)} \\ Hom_{C}(X, Γ Y) \ar[r]^{id} & Hom_{C}(X, Γ Y) \ar[r]^{ϕ_{X,Y}} & Hom_{C}(LX, Y)}

It remains to show that Im L = Im Γ. Suppose that X belongs to Im L. Then ηX is invertible, and therefore the map Hom_{C}(X, θY) is bijective for all Y. Using this fact for Y = X and Y = Γ X shows that θX is invertible. Thus Im L ⊆ Im Γ. The proof of the other inclusion is similar. □

Proposition A.4. For a localization functor L: C → C are equivalent: (1) The functor L admits a right adjoint.

(2) The inclusion functor G: Im L → C admits a right adjoint G_{ρ}.

(3) There exists a colocalization functor Γ: C → C such that Im Γ = Im L. In that case Γ ≅ GG_{ρ} and Γ is a right adjoint of L.

Proof. (1) ⇒ (3): See Lemma A.3.

(3) ⇒ (2): See Lemma A.1.

(2) ⇒ (1): Write L as composite L = GF as in Lemma A.1. Then the composite GG_{ρ} is a right adjoint of L, since G is a right adjoint of F. □

The left adjoint of a localization functor. We discuss the existence of a left adjoint of a localization functor.

Proposition A.5. For a localization functor L: C → C are equivalent: (1) The functor L admits a left adjoint.

(2) The functor F: C → Im L sending X in C to LX admits a left adjoint F_{λ}.

(3) There exists a colocalization functor Γ: C → C such that Γ ϕ is invertible if

and only Lϕ is invertible for each morphism ϕ in C. In that case Γ ≅ F_{λ}F and Γ is a left adjoint of L.

Proof. Denote by Σ be the class of morphisms in C that are inverted by L.

(1) ⇒ (2): Let L_{λ} be a left adjoint of L and write L = GF as in Lemma A.1. Then we have for X, Y in C

Hom_{C}(X, F Y) ≅ Hom_{C}(GX, GF Y) ≅ Hom_{C}(L_{λ}GX, Y) .

Thus L_{λ}G is a left adjoint of F.

(2) ⇒ (3): It follows from Lemma A.2 that F induces an equivalence C[Σ^{−1}] ⟶^{∼} Im L. Applying the dual assertion of this lemma shows that F_{λ} is fully faithful and that Γ = F_{λ}F is a colocalization functor. Moreover, the class of morphisms that are inverted by Γ coincides with the corresponding class for F, which equals Σ.

(3) ⇒ (1): Denote by Q: C → C[Σ^{−1}] the universal functor inverting Σ and let L̄, Γ̄: C[Σ^{−1}] → C be the induced functors satisfying L = L̄Q and Γ = Γ̄Q. It follows from Lemma A.2 that L̄ is a right adjoint of Q and that Γ̄ is a left adjoint of Q. Thus for X, Y in C we have

Hom_{C}(Γ̄QX, Y) ≅ Hom_{C}(QX, QY) ≅ Hom_{C}(X, L̄QY) .

It follows that Γ is a left adjoint of L. □

Remark A.6. Let C be a triangulated category and L: C → C an exact localization functor. Then statements (2) and (3) in Proposition A.5 admits the following equivalent reformulations:

(2′) The quotient functor C → C/ Ker L admits a left adjoint. (3′) There exists an exact colocalization functor Γ: C → C such that Ker Γ =

Ker L. The equivalence C/ Ker L ⟶^{∼} Im L has already been mentioned. The reformulation of (3) relies on the same argument: an exact functor inverts a morphism ϕ in C if and only if it annihilates the cone of ϕ.

References

1. D. J. Benson, Representations and Cohomology I: Basic representation theory of finite groups and associative algebras, Cambridge Studies in Advanced Mathematics, vol. 30, Cambridge University Press, 1991, reprinted in paperback, 1998.

2. D. J. Benson, Representations and Cohomology II: Cohomology of groups and modules, Cambridge Studies in Advanced Mathematics, vol. 31, Cambridge University Press, 1991, reprinted in paperback, 1998.

3. D. J. Benson, J. F. Carlson, and J. Rickard, Complexity and varieties for infinitely generated modules, II, Math. Proc. Camb. Phil. Soc. 120 (1996), 597–615.

4. D. J. Benson, S. B. Iyengar, and H. Krause, Local cohomology and support for triangulated categories, Ann. Sci. École Norm. Sup. (4) 41 (2008), 573–619.

5. D. J. Benson, S. B. Iyengar, and H. Krause, Stratifying modular representations of finite groups, preprint, arxiv:0810.1339.

6. D. J. Benson, S. B. Iyengar, and H. Krause, Stratifying triangulated categories, preprint, arXiv:0910.0642.

7. D. J. Benson and H. Krause, Complexes of injective kG-modules, Algebra & Number Theory 2 (2008), 1–30.

8. W. Bruns and J. Herzog, Cohen-Macaulay rings, Cambridge Studies in Advanced Mathematics, vol. 39, Cambridge University Press, Cambridge, 1998. Revised edition.

9. W. G. Dwyer and J. P. C. Greenlees, Complete modules and torsion modules, Amer. J. Math. 124 (2002), 199–220.

10. L. Evens, The cohomology ring of a finite group, Trans. Amer. Math. Soc. 101 (1961), 224– 239.

11. L. Fuchs, Infinite abelian groups. Vol. I, Pure and Applied Mathematics, vol. 36 Academic Press, New York-London, 1970.

12. P. Gabriel and M. Zisman, Calculus of fractions and homotopy theory, Ergebnisse der Mathematik und ihrer Grenzgebiete, vol. 35, Springer-Verlag, New York, 1967.

13. E. S. Golod, The cohomology ring of a finite p-group, Dokl. Akad. Nauk. SSSR 125 (1959), 703–706, (Russian).

14. J. P. C. Greenlees and J. P. May, Derived functors of I-adic completion and local homology, J. Algebra 149 (1992), 438–453.

15. M. Hovey, J. H. Palmieri, and N. P. Strickland, Axiomatic stable homotopy theory, Mem. Amer. Math. Soc., vol. 610, Amer. Math. Soc., (1997).

16. B. Keller, Deriving DG categories, Ann. Sci. École Norm. Sup. (4) 27 (1994), 63–102.

17. H. Krause, A Brown representability theorem via coherent functors, Topology 41 (2002), 853–861.

18. H. Krause, The stable derived category of a noetherian scheme, Compositio Math. 141 (2005), 1128–1162.

19. J. Lipman, Lectures on local cohomology and duality, Local cohomology and its applications (Guanajuato, Mexico), Lecture Notes. Pure Appl. Math., vol. 226, Marcel Dekker, New York (2002), 39–89.

20. A. Neeman, Triangulated categories, Annals of Math. Studies, vol. 148, Princeton Univ. Press, 2001.

21. A. Neeman, Colocalizing subcategories of D(R), J. Reine & Angew. Math., to appear.

22. D. Quillen, The spectrum of an equivariant cohomology ring. II, Ann. of Math. 94 (1971), 573–602.

23. A. S. Richardson, Co-localization, co-support and local homology, Rocky Mountain J. Math 36 (2006), 1679–1703.

24. B. B. Venkov, Cohomology algebras for some classifying spaces, Dokl. Akad. Nauk. SSSR 127 (1959), 943–944.

25. J.-L. Verdier, Des catégories dérivées des catégories abéliennes, Astérisque, vol. 239, Société Math. de France, 1996.

Dave Benson, Institute of Mathematics, University of Aberdeen, King’s College, Aberdeen AB24 3UE, Scotland U.K.

Srikanth B. Iyengar, Department of Mathematics, University of Nebraska, Lincoln, NE 68588, U.S.A.

Henning Krause, Fakultät für Mathematik, Universität Bielefeld, 33501 Bielefeld, Germany.